A Complete Bibliography of Publications in the ICES Journal of Marine Science (1960–1969)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.03

Title word cross-reference

14 [Gol68].

0-Group [Ile68].

170^{o} [HP69]. 1960 [Ruu61]. 1961 [Spä62].

2 [Bou63]. 2’-Diquinolyl [Bou63]. 2’-Diquinolyle [Bou63].

40 [Thr69].

Âge [eJPT66].

A. [Gra64]. Abbau [Sze68a]. Abnormal [dV69]. Absolute [Gol68]. Abundance [Dow64, PA62, Zij67]. acanthias [HM64, Hol66]. Acclimatization [Bow66, EZ63]. Acoustic [Has62]. Action [GC68].

Activities [Sor64a]. Activity [Gol68, RKT67]. Adapting [NP64]. Adriatic [Vat62]. Adsorption [TH68]. Adult [Bow66, Rai61b]. aeglefinus [Liv62]. Affecting [LF69]. Africa [Lon60, Lon64]. Africains [eJPT66]. African [eJPT66, Lon65a]. Age [Alb68, Chr64, Dut69, GM64, KQ68, Pan62, Pan63, Pan69, RS67, Smi68, eJPT66]. Age-Structures [RS67]. Aid [Bac66, Hie62, SL62]. Alf [Ruu61]. Alfax [Jon69]. Alfredo [Can61]. Algal [Har68]. Alkalinity [Gri60]. Alm [Lin62]. Alternative [Cus62a]. Ammonia [Bee64, Vac62]. among [Dow64]. Analysis [Gul66a, IJ62]. anchoita [dC66]. Anchovy [dC66, dM69]. Andersson [Hög69b]. Andreas [Hög69b]. Anglian [Bol62, Bur63, Cus68b, Ile67, Ile68]. Animals [Hei62, Mar63, Ree64]. Annual [Bra63, HE67]. Annuel [Bra63]. Antarctic [Gul66b, Gul68b]. Anton [Spä62]. Anwendung [Sze68a]. any [Mor68]. Apercu [GC68]. Apparatus [Fah64, Lov60]. Apparent [Jon60]. Applicable [RW68, San69]. Application [Sze69a, Sze69b, Sze68a]. Aquaria [Ryl66]. Aquatic [Ano60a]. Area [Mar60, VC66]. Areas [SF68, Wil69b]. Argentina [Bol64]. Argentine [dC66]. Arni [Ola67]. Arrival [Aas62]. Artemia [Mit63, Ril66b, SV69]. Arthur [Cul68]. Artificial [Sim64]. Arvid [Hög69a]. Asdic [SL62]. Assessment [BB65]. Association [Dow64]. Atlantic [CR61, HB67, Lin63]. Attraction [HM68]. Aufund [Sze68a]. auna [CC66]. Automatic [MW61]. Autumn [MP60, Zij69]. Autumn-Spawning [Zij69]. Avoidance [BP65].

B [TH68, WK67]. Back [Kip62]. Back-Calculation [Kip62]. Bacteria [CW65, MS65]. Bacterial [Sor64a]. Baltic [Gri60]. Bank [Mar60]. Barium [MT66]. Barnacle [Moy63]. Basin [Fur68]. Bassin [Fur68]. Bathyrheograph [BFRT67]. Bay [SF68, Bra63]. Behaviour [Har68, SF68, Str61, dG69]. Benthic [MP61]. Bermuda [HRG60, Men60, MR61]. Berried [SF68]. between [Aas62, AL63, BeB60, Dow64, Paa60, RS67]. Bibliography [Ano60a]. Bight [VC66]. Biochemical [Lin65, RAL64, RKT67]. Biodeposits [HMA66]. Biological [SJI68]. Biology [Lon65a, dFT67]. Biomass [GC68]. Biomasse [GC68]. Biometric [PT67]. Bionomics [Lon64]. Bivalves [CA64]. Black [Sor64a]. Blkr [eJPT66]. Bløden [Rai61b]. Blood [Wil67a]. Bloom [Hal60]. Bluefin [Mat62]. Body [MP60, Lin63]. Body/ [Lin63]. Bomb [Bir69]. Bonga [Bai61, Bai63]. Books [Ano66a, Ano67a, Ano67b, Ano68a, Ano68b, Ano68c, Ano69a]. Bottles [PNYR64]. Bottom [CC66, Day60, Las65, Mac67]. Bottom-Plankton [Mac67]. breakdown [Sze68a]. Breeding [Mil64]. Britain [Bow66, Mit63, RT63, Ril66b, SRT63, She63a, She63b]. Brown [Kip62]. Browns [Day60]. Bruun [Spä62]. Bubble [CW65]. Buch [Pan62]. Buchan [Cra60]. Bucket [Alb69]. build [Sze68a]. build-up [Sze68a]. Buoyancy [SLI65]. Burning [Chr64].

C [Gol68]. C. [Lar62, eJPT66]. Caecae [Rai62]. caerulea [LT62]. Calcium [MT66, Sim64]. Calculate [Sze69a]. Calculation [Kip62]. Calculations [Has62]. Calorific [Sit67]. Calorimetry [Bir69]. Cancer [HE67]. Cap [BeB60]. Cape [BeB60]. Carapace [Hep66]. Carbon [Hop68, Sze69a, Sze69b, WS67, Sze68b]. Carbon-14 [WS67]. Carbonate [Sim64]. Carbone [Sze69b]. Carried [TH68]. Catch [All67, Bol62, TF65]. Catches [Cra60, Gul61, Gul64a, Gul66b]. Catfish [Dut69]. catodon [Roe69]. Caught [Fin65]. Cells [Har68]. Central [Jon62a, Sor64b]. Century [Cus68b]. Certain [RW68, San69]. Certains [RW68]. Chaetognath [Bee64, RS67]. Chaetognatha [RS68]. Chamber [Tun67]. Changes [All67, Gul61, Gul64a, Gul68c, SCC68, Zar60]. Characteristics [Gar61, IJ62, Jon69, PNYR64]. Characters [HB61]. Chemical [Bla69, Sze69a, Sze68a]. Chemisches [Sze68a]. chez [eJPT66]. Chlorinity [Mor68]. Chloroform [Jon63]. Chlorophyll [HHLHS65]. Chukchi [PA62]. Circulating [Fah64]. Circulation [RT63, SRT63, She63a, She63b]. Citations [Ano66e]. Clam [Ans68]. Clarias [GH61]. Class [IJ62]. Clogging [SCC68]. Closed [SRT63, RT63]. Clupea [BH63, HB61, HB67, IJ62, Jon62b, Kab63, Nie60b]. Coalfish [Mar60]. Coast [Roe69, Hol60a, Tho64]. Coccidioses [Kab63]. Cockle [Tay60a]. Cod [Bla69, Hol60a, Hol60b, Liv62, Lov62, Mar60, SLI65, Tay60b, Wil67a]. Cod-End [Liv62]. Colonies [AL63]. Commercially [Bar68]. Communities [Hei62, Hei69]. Comparaison [Bou63]. Comparative [GM69, Sze68b]. Comparées [GM69]. Comparison [Dow64, Moy63, Bou63]. Competition [Rot67]. Complication [Nie64]. Composition [Bla69, RAL64, TF65]. Conant [Bee64]. Concentration [Bis62b, CW65]. Concentrations [Bis62a]. Concept [Gul68a]. Condition [AL63, Lov62]. Conditions [Bow66]. conical [GM69]. Coniques [GM69]. Conseil [Ano66c]. Conservation [GC68]. Considerations [LE65]. Consumption [Bar68, LT62, Sze68b]. Content [Hop68, Jon63, Lin65, Sze68b]. Continental [BeB60, BeB60]. Continuous [BFRT67]. Continuously [Har68]. Controlled [Fah64]. Cooperative [J.62]. Copepod [RS67]. Copepodfeeding [Har68]. Copepods [TH62]. Copepodsin [Cus68c]. Copper [Bou63]. Cores [Cra65]. Correlation [Aas62, Hie62, IJ62]. Coulter [CNF68]. Council [Ano65a, Sme68]. Counter [CNF68]. Counting [MS65, MW61, Mit63, WS67]. Countries [Ano65a]. courants [HP69]. Course [NP64]. Cox [Cul68]. Crab [HE67]. Crawfish [Hep66]. Critical [Cus62a]. Crop [Jon62a]. Crustacean [Hop68]. Cuivre [Bou63]. Culture [Bow66, MS65, Mit63, RT63, Ril66b, SRT63, She63a, She63b]. Cultured [MS68]. Culturing [SP66]. Current [BFRT67, Jon62b, Ano60a, HP69]. Currents [BFRT67, Bis60a]. Curve [Rif60]. Curves [Far60a]. Cuthbert [Car69]. Cuvier [Bai61, Bai63]. CV [eJPT66]. Cycle [Bra63, CR61, Mil64]. Cycles [Hei62]. cynoglossus [Bow60]. Cysts [SV69].

Dannevig [Ruu61]. Dans [Sze69b, HP69]. Dark [PNYR64]. Data [Sme63]. December [Spä62]. Decision [Pat64a]. Deep [Bac66]. dehydrogenase [RKT67]. Demersal [Pea62]. Demkaliumpermanganatverbrauch [Sze68b]. dempotassium [Sze68b]. Density [She68]. Dependence [Cra60]. Depth [BFRT67, Cus62a, VC66]. Depth-Current [BFRT67]. Desbrosses [Fur64]. Describingthe [Har68]. Deseado [Bol64]. Design [HG66]. Dessen [Sze68a]. Detection [Has62]. Determination [Bou63, Chr64, CNF68, GM64, HHLHS65, Mor68, Pan62, RW68, SLI65, Sze69b, RW68, eJPT66, eJPT66]. Determining [CA64, Tra60]. Detritus [Kre61]. deux [eJPT66]. Developing [Fah64]. Development [Ryl66, dC66]. Device [MP61, Mit63]. Devices [Boo61]. Diatoms [Moy63]. Diethyldithiocarbamate [Bou63, Bou63]. Different [Bis62a, Far60a, GM69]. Différentes [GM69]. Differentials [Gul64c]. Differing [IJ62]. Digestive [dG69]. Digital [TM65]. dijficilis [Hop68]. Dimensions [GM69, Has62]. Diquinolyl [Bou63]. Diquinolyle [Bou63]. Direct [CA64, MS65]. Directly [San69]. Discovery [JM62]. Discrimination [Gul66a]. Displaced [SF68]. Distributed [Har68]. Distribution [Bai63, BB68, BS69, Bis60a, Bis61a, Bis61b, CHT61, JH63, MCM63, Pea62, PA62, Sor60, VR60, VC66, HP69, HP69]. Diurnal [VC66]. Diving [Str61]. Dobson [Gra64]. Dolioles [Bra63]. D’Organismes [GC68]. dorsalis [Bai61, Bai63]. Dosage [Bou63, Sze69b]. Downs [Cus68a]. Drift [Bol62]. Drifter [Bur63]. Drifters [HG66]. Dry [SV69]. Due [SCC68, Nie64]. Dunmore [BB65, WF66]. Duration [Ile64]. During [Cus68a, SP66, She63a, She63b]. Dutch [ZdV63]. Dyeing [Alb68].

Early [Bai61]. East [Bol62, Bur63, Cus68b, Ile67, Ile68, MPS60]. Eastern [CR61, JH63, Lon65b, PA62]. Eau [Bou63, Sze69b]. Echo [Has62, Rai64, VC62, VC66]. Echo-Strength [Has62]. Echo-Traces [Rai64]. Echoes [MW61]. Ecological [BS69]. Edible [Bar68, HE67]. edulis [WK67, MS68, She68]. Effect [All67, Bis60a, Bis61a, Bis61b, Bow66, Far60a, Gul61, Gul66b, Jon62a, Jon63, MS68, She68, VC62]. Effectiveness [Bar64, CHT61]. Effects [Gul64a, Ril66b, Sim64]. Efficiences [GM69]. Efficiencies [WS67, GM69]. Efficiency [Jør60, SCC68]. Egg [BH63, HB67]. Eggs [Bow66, Fah64, Hie62, LT62, Pea62, SLI65, dC66]. Egypt [EZ63]. Eighteenth [Cus68b]. Einarsson [Ber68]. Electronic [Gri68, JM62, Mit63]. Eliminating [Gol68]. End [Liv62]. Engraulis [dC66]. entre [BeB60]. Epinephelus [Men60]. Equations [Tay62]. Equatorial [HP69]. équatoriaux [HP69]. Equipment [Fin65, Gul66a]. Erin [She63a, She63b]. Errors [Gol68]. esculenta [Gar61]. Espagne [BeB60]. Espèces [eJPT66]. Essai [Fur68]. Essay [Coo63, Fur68]. Estimates [Far60a, Zij67]. Estimating [All67, Cus62a, Rif60, San69]. Estimation [Gul61, Han65, HE67, KQ68, Spe64]. Estuary [Bur69]. Ethmalosa [Bai61, Bai63]. Étude [Bra63, RW68]. Études [GM69]. Euphausiid [Hop68]. European [WK67]. Evaluation [Min68]. Evidence [Hol60a, Jon62a]. Examination [Smi68]. Excretion [Bee64]. Experiment [She63a]. Experimental [AL63, HB61, PNYR64]. Experiments [HM68, MJ67, MS65, Nie60b, She63b, WK67]. Exploitation [KQ68]. Exploited [Gul68a, San69]. Exploration [Ano65a, Sme68]. Extensive [VC62].

Facilitate [Has62]. Factor [BR60]. Factors [Gul64c, LF69, Lon60, dG69]. Fage [Ber65]. Farms [Vat62]. Faroese [Hol60a]. Fauna [Vat62]. Fecundity [HM64, Pan63, RH67, Tho64]. Feeding [Bai63, Ril66b, Roe69, dG69]. Female [SF68]. Fertility [MPS60]. Field [HM68]. Filets [GM69]. Filtering [SCC68]. Filtration [Alb69]. Fish [Ano65a, Bar68, Bis60a, Bis61a, Bis61b, Bow66, BR60, Chr64, Fah64, Far60a, FAR60b, Gri68, Gul68a, Has62, Hie62, HM68, JM62, KQ68, Lon65b, Men60, MW61, Mit63, Ols64, RT63, Ril66a, Ril66b, SRT63, She63a, She63b, Vat62, Wil67b]. Fisheries [Ano60a, Rai61b]. Fishery [Bol62, Bur63, Cra60, Cus68b, Gul68c, Lon60, Rot67, Rou62, Sor64c]. Fishes [Cus69, Lon65a]. Fishing [ZdV63]. Flagellates [Moy63]. Flat [WK67]. Flatfish [dG69]. Fleets [Bur63]. Floating [HM68]. Flounders [PA62]. Flowmeter [TM65]. Flume [MJ67]. Fluorescent [HMA66]. Fluorometric [HHLHS65]. Focus [BeB60]. Following [JM62]. Food [Bai63, Hol66, Mar60, Men60, Moy63, Ric63, Roe69, dM69]. Foraminifera [Bol64]. Form [Lov60]. Formalin [AW62]. Fractions [Bar68]. Frederik [Spä62]. French [BeB60, Bou63, Bra63, Fur68, GC68, GM69, HP69, RW68, Sze69b, eJPT66]. Fresh [Hop68]. Freshwater [Sit67]. Frioriksson [Ola67]. Further [Jon62b].

Gadus [Bla69, Hol60a, Rai62, Rou62, SLI65, Wil67a]. Gear [Gul61, Gul64a, Rot67]. Gehalt [Sze68b]. Geiger [WS67]. General [GM69]. Générales [GM69]. Geographic [Ano66b]. Geographical [Ans68]. Georges [Mar60]. German [Sze68a, Sze68b]. Gill [Gar61, GH61]. Girard [LT62]. Glyptocephalus [Bow60]. Golfe [Sze69b]. Gompertz [Rif60]. Gonadial [Mil64]. Grab [Las65]. Graham [Gar61]. Granqvist [Lis66]. Graphical [Han65]. Grazing [Cus68c]. Greenland [Nie60b]. Greff [RS68]. Ground [SL62, Rai61b]. Group [RW68, Ile67, Ile68]. Groupes [RW68]. Groups [IJ62]. Growth [AL63, Ans64, Ans68, Bow60, Dut69, Far60a, Han65, HE67, Hol60b, IJ62, Ile67, Ile68, Jon60, Kip62, MS68, Pan62, Pan63, Pan69, PA62, Rif60, She68, Tay60a, Tay60b, Tay62]. Growth-Rate [Hol60b, Tay60b]. Guide [Ano65a]. Guinea [Lon65b]. Gulf [Lon65b, CHT61, Mar60, Sze69b]. Gunnar [Lin62, Lis66]. guttatus [Men60].

Habits [Bai63, Mar60, Roe69, Zar60]. Hadddock [MCM63]. Haddock [Jon60, Liv62, Mar60]. Hamilton [Pan62]. Hard [Ans68]. harengus [BH63, HB61, HB67, Jon62b, Kab63, Nie60b]. Hawaiian [Str61]. Headless [Wil67b]. Heat [Bir69, IS69]. Helen [Fra60]. helgolandica [RS68]. Henrik [Hal61]. Herbivorous [Cus68c]. Hermann [Ber68]. Herring [Aas62, BH63, BP65, Bol62, Bur63, BB65, CHT61, Cra60, Cus68a, Cus68b, Dow62, HB61, HB67, Ile64, Ile67, Ile68, Jon62b, Kab63, MP60, Nie60b, PZD65, Rai61a, Ric63, Ric60a, Sym64, WF66, Zij67, Zij69]. High [She68, She63b, Vat61]. hispida [Bee64]. Histories [Hei62]. History [Bai61, Wil69a]. Hodgson [Car69]. Hubbs [dC66]. Hydrodynamics [MJ67]. Hydrogen [Bis62a]. Hydrographique [Sme68].

I-group [Ile67]. Iceland [Roe69]. ICES [Sme63]. Identifying [Hie62]. II [JM62, Ile68, Lin65, She63a]. III [RT63]. im [Kre61, Sze68b]. Immature [PZD65]. Implements [FAR60b]. Importance [BP65]. Important [Bar68]. Improved [Lov60]. Improvement [Las65]. Incidence [Kab63]. Incorporated [WS67]. Index [Ano60b, Ano61, Ano62, Ano64, Ano65b, Ano66b, Ano66c, Ano66d, Ano66f, Ano66g, Ano68d, Ano69b]. India [Bur69]. Indication [RS67]. Inductive [Ger60]. Inexpensive [TM65]. Influence [BH63, Bol62, Nie64, Sor64c]. Inhibitory [Sim64]. Initial [San69]. Inorganic [Bee64, Jon63]. Inshore [Rou62, TH68]. integer [RAL64, RKT67]. Intensities [NP64]. Interim [Gul64a]. International [Ano65a, Sme68]. intestinalis [Wil69a, Wil69b]. Inverted [Lov60]. Investigations [Rai64, SL62, RKT67]. Ion [Bis62a]. Ireland [BeB60]. Irish [JH63, Bow60, MT66, Ric63]. Irminger [Ste68]. Island [Jon62a, SF68]. Islands [Jon62a, Hol60a]. Isle [She63a, She63b]. IV [RKT67, She63b].

Jar [SP66]. Järvi [Hal61]. Journal [Ano66c].

Karl [Hög69b]. Kohlenstoff [Sze68b].

L [AL63, Ans68, MS68, RT63, SV69, WK67]. L. [Ans64, Bla69, BH63, Bow60, GM64, HE67, HB61, HB67, Hol60a, HM64, Hol66, IJ62, Jon62b, Kab63, Kip62, Lin63, Mit63, Nie60b, PT67, RH67, Ril66b, Roe69, Rou62, Ryl66, Sor64c, SLI65, Tho64, TF65, Wil67a, dFT67, dV67, dV69]. Laboratory [SP66]. Lagoon [Vat61]. Lake [EZ63]. Lamellibranchs [Jør60]. Large [Mat62]. Larvae [Bis60a, BH63, Bow66, Fah64, LT62, MS68, Moy63, Ryl66, dC66]. Larval [Bis60b, Bis61a, Bis61b, Bis62b, Bis62a, CHT61, Far60a, Mar60, MCM63, Rai64, Ril66b]. Latex [Ril66a]. Layer [VC62]. Lecture [eJPT66]. Ledge [Day60]. Length [Hep66]. Lethal [Bis60b]. Levels [Ril66b]. Life [Bai61, Hei62, Wil69a]. Light [BP65, Nie64, NP64, PNYR64]. Limit [All67]. Line [Fin65]. Linn [EZ63, Pan63]. Linn. [Dut69]. Linnaeus [Liv62]. Lion [Sze69b]. Lipid [Lin65]. Liquid [Ril66a, WS67]. l’Irlande [BeB60]. Lists [Sme63]. Living [Bol64]. Lobster [TF65, dFT67]. Lobsters [SF68, Tho64]. Loch [Mil64]. Louis [Ber65]. Low [Bis62b, NP64]. Lowestoft [RT63, Rou62, SRT63].

M [Can61, Thr69]. Mêmes [GM69]. Macroplanctoniques [GC68]. Macroplankton [GC68]. Mailles [GM69]. Maine [CHT61, Mar60]. Man [She63a, She63b]. Management [Gul68b]. Manx [Sym64]. Marginal [Gul68a]. Marine [BeB60, Bow66, Cra65, CNF68, Dow64, HMA66, IS69, Lin65, Mar63, MT66, MS65, Mit63, Nie64, NP64, RAL64, RKT67, RT63, Ril66b, SP66, SRT63, She63a, She63b, VR60, WS67]. Marini [dC66]. marinus [RH67, Sor64c]. Marking [Ril66a]. Marks [Ano65a]. Marquesas [Jon62a]. Massachusetts [Day60]. Masses [RS67]. material [Sze68a]. Materials [Sze68a, TH68]. Mathematical [Har68]. Maturation [Ile64]. Mature [Ans64]. Means [BFRT67, GM64, Smi68]. Measured [BFRT67]. Measurement [Gol68, Has62, Hep66, Lov62, TH62]. Measurements [Ger60, WS67]. Measuring [Boo61, Wil67b]. Mechanical [IS69, TM65]. Mediterranean [Fur68, Hal60]. Mediterranée [Fur68]. Meere [Kre61]. Meerwasser [Sze68b]. Meiof [CC66]. Melanogrammus [Liv62]. Members [Ano65a]. mentella [Jon69, Sor64c]. Mer [Bou63, Sze69b, Bra63]. Mercenaria [Ans68, AL63, Ans64]. Meristic [HB61]. Merlangius [GM64]. merlangus [GM64, Rou62]. Mesh [Gul64c, Jon60, Lon60]. meshes [GM69]. Metabolic [Tay62]. Metabolites [Sim64]. Metamorphosed [She63b]. Meter [Ger60]. Method [CA64, Cus62a, Fin65, MW61, Rif60, San69, Tra60, Wil67b]. Methods [All67, Bou63, CHT61, Bou63]. Microbalance [Ree64]. Microflora [Sor64b]. Micronectonic [GC68]. Micronectoniques [GC68]. Micropipette [Thr69]. Microplankton [Alb69]. Microscope [Lov60]. Microscopes [Thr69]. Microzooplankton [BS69]. Mid [Rai64]. Mid-Oceanic [Rai64]. Midwater [VC62]. Migration [BP65, Mat62, Pat64a, SS63]. Migrations [Hol60a, ZDG64]. Migratrion [Pat64b]. militaris [Dut69, Pan63]. Mise [BeB60]. Model [Har68, Sze69a, Sze68a]. Modell [Sze68a]. Modification [Bir69, HB61]. Modulation [CJG66]. Molander [Hög69a]. Mollusc [San69]. Molybdenum [Bur69]. morhua [Bla69, Hol60a, SLI65, Wil67a]. Morphologie [BeB60]. Morphology [BeB60]. Mortality [San69, She68, Tay60a]. mossambicus [GH61]. Movement [SS63]. Movements [Jon62b, JM62, SF68]. Multi [BT67]. Multi-Purpose [BT67]. Multiple [Rot67]. Multiple-Species [Rot67]. Mysidacea [Lin65]. Mytilicola [Wil69a, Wil69b].

Narragansett [SF68]. Natural [San69, Sim64]. Nauplii [Mit63, Ril66b]. near [Jon62a]. Nematoscelis [Hop68]. Neomysis [RAL64, RKT67]. Nephrops [PT67, Tho64, TF65, dFT67]. Net [Bar64]. Nets [BP65, Bol62, Gar61, GH61, MJ67, SCC68, GM69]. Nicolai [Ruu61]. Nile [Hal60]. Nitrate [Ste68]. Nitrate-Phosphate [Ste68]. Nitrite [VR60, HP69]. Nitrogen [Hop68, Sit67]. Non [Ano65a]. Non-participant [Ano65a]. North [LF69, Rai61a, VC66, Zij67, CR61, Gul68c, Ile67, Ile68, Lov62, MPS60, MP60, PZD65, Vat62, Zij67, Zij69]. North-East [MPS60]. North-Eastern [CR61]. North-Western [Zij67]. norvegicus [PT67, Tho64, TF65, dFT67]. Norway [Hol60a, Tho64, TF65, dFT67]. Norwegian [Aas62]. Note [AW62, Cra60, Gul64a, HG66, Min68, Rai62]. Notes [Mar63]. Numbers [Ric60a]. Nylon [Gar61, GH61].

Objects [HM68]. Observations [AL63, BS69, Hal60, Jon62b, Liv62, LE65, MP60, PT67, Rai64, Ric60a, Rou62, Ryl66]. Observing [SY60]. obtained [BeB60]. obtenus [BeB60]. Occidental [Fur68]. Occurrence [Rai61b]. Occurrences [Bol64]. ocean [HP69, HP69, MPS60, Sor64b]. Oceanic [Rai64]. Oceanographic [LE65, SJI68, Sme63]. Oceans [Hei62]. off [Day60, HRG60, Lon60, MR61, Roe69, Rou62]. Offenen [Sze68a]. Offshore [SF68]. Ogilvie [Fra60]. Open [She63a, She63b, Sze68a]. Optimum [KQ68]. Order [Pea62]. Organic [MR61, Sor60, Sze69a, Sze68a, Sze68b]. Organischem [Sze68b]. Organischen [Sze68a]. Organisms [Bar64, GC68, MT66, RW68, SP66, Sit67]. Orientation [SS63]. Origin [Pea62]. Ortegal [BeB60]. Oslofjord [BFRT67]. Osmoregulation [LT62]. Osteogeneiosus [Pan63]. Osteogneiosus [Dut69]. Ostrea [MS68, She68, WK67]. Other [Chr64]. Otolith [Alb68, Min68, Rai61a, Rai61b, Smi68]. Otolithes [eJPT66]. Otoliths [Bla69, Chr64, GM64, eJPT66, WF66]. Otterstrøm [Lar62]. Ouest [eJPT66]. Ouest-Africains [eJPT66]. Outputs [Bir69]. Overview [GC68]. Oxidation [Vac62]. Oxygen [Bis62b, LT62]. Oyster [HMA66, MS68]. Oysters [Mil64, She68, WK67].

P [MPS60]. Pacific [MPS60, Sor64b, Hei69, HP69, Jon62a, LT62, Tay60a]. Pacifique [HP69]. pagurus [HE67]. Palinurus [Hep66]. Pangasius [Pan62]. Parameters [Ans64, Han65, Rif60, Tay62]. participant [Ano65a]. Particle [Jør60]. Particles [HMA66]. Particulaire [Sze69b]. Particulate [Sze69b]. Patagonia [Bol64]. Pectoral [Pan62]. Pelagic [Hie62, HM68, Jon69, SP66]. Period [Cus68a]. permanganate [Sze68b]. Peruvian [dM69]. Peters [GH61]. Phaeocystis [JH63]. Phenomenon [dV67]. Phosphate [Jon63, Ste68]. Phosphorus [Bee64]. Physeter [Roe69]. Physical [Wil69b]. Phytoplankton [Hal60, HRG60, NP64, Paa60, Sor60, Spe64, VR60, WS67]. Pierre [Fur64]. Pigmentation [dV69]. Pigments [Spe64]. Pinger [Bac66]. Plaice [Bow66, Gul68c, Ryl66, She63a, She63b, dV69, RT63, Ril66b, SRT63]. Plancton [GM69]. Planctoniques [RW68]. Plankton [AW62, BT67, CR61, Hei62, Hei69, Hie62, Mac67, MJ67, SCC68, Thr69, GM69]. Planktonic [Bee64, RW68]. Plateau [BeB60]. platessa [RT63, Ril66b, Ryl66, dV69]. Pleuronectes [RT63, Ril66b, Ryl66, dV69]. Pleuronectiformes [Pea62, dG69]. point [BeB60]. Pole [Fin65]. Polymorphism [Wil67a]. Polynemid [Lon65a]. Pomeranian [ZDG64]. Pomolobus [Dow62]. Pond [Bow66]. Population [BB65, RS67, She68, Sor64c]. Populationof [Har68]. Populations [KQ68, San69, dV69]. Port [She63a, She63b]. Possible [BR60, Har68, dV69]. Post [Ril66b]. Post-larval [Ril66b]. Potential [KQ68]. pour [Bou63]. poutassou [Rai62]. Power [IS69, ZdV63]. Pre [Cra60]. Pre-Spawning [Cra60]. Precipitation [Sim64]. Predation [BR60, RS68]. Preliminary [AL63, Lin63, Nie60b]. Presence [VC62]. Preservation [AW62, Mar63]. Preserved [Hop68]. Pressure [Bis61a]. Primary [CNF68, Gol68, MPS60, Paa60, Sor60, Sor64a, Vat61]. Problem [Min68].

Procedes [GC68]. Process [Pat64a]. Processes [GC68]. Production [MR61, Paa60, Sor60, Sor64a, Vat61]. Productivity [CNF68, Gol68, IS69, MPS60, Nie64, WS67]. Profile [Rum61]. Progress [SY60]. Proportioning [Gri68]. Proportions [MP60]. Proteins [Wil67a]. pseudoharengus [Dow62]. Pseudotolithus [eJPT66]. Puerto [Bol64]. Pumping [CA64]. Purpose [BT67]. Pyloric [Rai62].

Qualitative [Min68]. Quality [MS68]. Quantitative [CC66, Sor64b]. Quantity [VC66]. Quarun [EZ63]. Quick [All67].

R [Hög69a]. R.R.S. [JM62]. Racial [Sym64, Zij69]. Rade [Bra63]. Ragnar [Kor67]. Ramalho [Can61]. Random [SS63]. Randomly [Har68]. Range [Ans68]. Rate [Ans68, CA64, Hol60b, Jør60, Tay60b]. Rational [Pat64a]. Reactions [Bis62b, Bis62a]. Reading [Alb68, Smi68, eJPT66]. Rearing [RT63, SRT63, She63a]. Received [Ano66a, Ano67a, Ano68a, Ano68b, Ano68c, Ano69a]. Recevied [Ano67b]. Recorder [BFRT67]. Recording [Rum61]. Redfish [Jon69, Rai64, RH67, Sor64c]. Reducing [BR60]. Reef [Men60]. reference [RKT67]. References [Ano66e]. Regularity [Cus69]. Regulation [Gul66b]. Related [Bar64]. Relation [Jon62b, MR61, dG69]. Relationship [AL63, Lin63, Paa60]. Relationships [Ste68]. Relative [Dow64, ZdV63]. Relativelydense [Har68]. Reply [Pan69]. Report [J.62, Lin63]. Research [J.62, SJI68]. Resources [Gul68b, Lon65b]. Respiration [SV69, TH62]. résultats [BeB60]. Results [Nie60b, ZDG64, BeB60]. Retention [Jør60]. Review [Bal62, Bir61, Coo68, Coo63, Cus62b, Eng62, Fra64, Gul64b, Pre63, Ric60b]. Reviews [Ber60, Bic67, Cha69, Cus61, Cus67, Kor69, Nie60a, Sai69, Sme61, Str60a, Str60b]. Rhodamine [TH68, WK67]. Rhodamine-B [TH68]. Rhode [SF68]. Risso [Rai62]. River [Dow62]. Rivers [ZDG64]. Roland [Cul68]. Running [SP66]. Ryan [Mil64].

Sagitta [Bee64]. salar [Lin63]. salina [Mit63, Ril66b]. Salinity [Bis61b, Ger60, Mor68, She63b]. Salmo [Kip62, Lin63]. Salmon [Lin63, Pat64a, Pat64b]. Salmonids [Bis60b, Bis62b, Bis62a]. Salmono [SS63]. Salpes [Bra63]. Salps [Bra63]. Salt [Fah64, Vat62]. Salt-Water [Fah64, Vat62]. same [GM69]. Sampler [Bar64, BT67, CC66, Cra65, Mac67]. Samplers [Bar64]. Samples [AW62, Hie62, MP61]. Sampling [CHT61]. Sardine [LT62]. Sardinops [LT62]. Sargasso [HRG60, MR61]. sauna [SV69]. Scale [Gri68, Lin63]. Scales [Kip62]. Scanning [CJG66, Gul66a, JM62]. Scavenging [CW65]. Schooling [BR60, Ols64]. Sciaenidae [Lon64, eJPT66]. Sciaenidés [eJPT66]. Sciences [Ano60a]. Scintillation [WS67]. Scotland [Mil64, TF65]. Scottish [Kab63, Rai64, Tho64]. Screening [MP61]. SDDC [Bou63]. Sea [Ano65a, Boo61, CW65, CJG66, Cus68c, Dow64, HG66, JM62, Jon63, MP61, She63a, Sim64, Sor60, SY60, Sze69b, TH62, VR60, Vac62, Zar60, ZDG64, Sze68a, Bow60, CR61, Gul68c, HRG60, Ile67, Ile68, JH63, LF69, Lov62, MT66, MR61, MP60, PZD65, PA62, Rai61a, Ric63, Sme68, Sor64a, Ste68, VC66, Zij67, Zij69]. Sea-Surface [HG66]. Season [Cus69]. Seasonal [Bol64, CR61, Hei62, TF65]. Seawater [Bou63, Mor68, Sze69a, Sze68b]. Sebastes [Jon69, RH67, Sor64c]. second [Ile67]. Sector [CJG66, Gul66a, JM62]. Sector-Scanning [JM62]. Sedimentation [Lov60, Tun67]. Sediments [HMA66]. See [Sze68a]. Selected [Dow64]. Selection [Bur63, Gar61, GH61, Gul64c, Jon60, Lon60]. Selectivity [Gul61, Gul64a]. senegalensis [eJPT66]. Sensorial [dG69]. Separate [IJ62]. Separation [MP61]. Serious [Gol68]. Service [Sme68]. Shallow [Rum61]. Shelf [BeB60]. Shellfish [Bar68]. Shipboard [SP66]. Shoaling [BP65]. Shoals [Jon62b, JM62]. Short [Cra65]. Silicate [BB68]. Similarity [RS67]. Simple [MP61, Ree64, Thr69]. Single [LT62, TH62]. Siphonate [CA64]. situ [Ger60]. Size [All67, Bar64, BH63, Ric60a, RW68, dC66]. Skipjack [Str61]. Slope [BeB60]. Small [Bir69, Ree64, Ril66a, SP66, SJI68]. Sodium [Bou63]. Solea [EZ63, dV67]. Soles [Chr64, dV67]. Soluble [Jon63]. Some [All67, Ans64, BS69, Cus69, Jon69, LE65, Mar63, MP60, PNYR64, PT67, Rai61b, Rou62, Ano65a, Bol64, Lin65, Rai64, Sim64]. Sømme [Ros62]. Son [Sze69b]. Sonar [CJG66, JM62]. Sous [BeB60]. Sous-Marine [BeB60]. South [BeB60, PA62]. South-Eastern [PA62]. Southampton [BB68]. Southern [Bur69, Day60, LF69, Rai61a, VC66]. Spain [BeB60]. Spärck [Kor67]. Spawning [Aas62, AL63, Cra60, Cus69, MP60, Pan63, SL62, Tho64, Zar60, Zij69]. special [RKT67]. Species [Bar68, Dow64, Rot67, eJPT66]. Speed [Bar64, Dow62]. Sperm [Roe69]. Spines [Pan62]. Sprat [IJ62]. sprattus [IJ62]. Spurdog [HM64, Hol66]. Squalus [HM64, Hol66]. Stages [Ile64]. Standard [Bir69]. Standing [Jon62a, Paa60]. Station [Jon69, MPS60]. Steuer [Wil69a]. Stimulation [IS69]. Stock [BB65, Cus68a, Paa60, WF66]. Stocks [Gul68a, Sym64]. Stratified [BFRT67]. Streams [VC62]. Strength [Has62]. Strontium [MT66]. Structure [Smi68, Sor64c, Zij69]. Structures [RS67]. Studies [BB65, Dut69, GM69, Ile67, Ile68, Lin65, Pan63, Rai61a, RAL64, RKT67, SP66, Sym64, WF66, Sze68b]. Study [CHT61, MCM63, RW68, Sor64b, dV69, Bra63]. Subject [Ano66f]. Submarine [SJI68]. Subsequent [Rai61b]. Substandard [Mor68]. succinic [RKT67]. Sud [BeB60]. Surface [HG66, Smi68, dV67]. Survey [Lon65b]. Survival [Bis60a, Bis61a, Bis61b, Far60a, Ols64]. Survivals [She63b]. Suspension [TH68]. Sven [Ros62]. Swimming [Bar64, Dow62, dV67]. Synthèse [Fur68]. Synthesis [Fur68]. System [Gri68, HP69, Rum61, dG69]. Système [HP69].

T [Gra64]. Table [Hie62, IJ62]. Tagging [Fin65, FAR60b, Nie60b, San69, ZDG64]. Tailles [RW68]. Taken [Ric60a, Roe69]. Taking [Cra65]. Talus [BeB60]. Target [Gul66a].

Taxonomic [Ano66g]. Technique [Alb68, Chr64, Fin65, MS65, Ril66a, RW68, TH62]. Television [Liv62]. Temperature [Boo61, Cra60, Day60, Dow64, Fah64, Hol60b, Rum61, Tay60a, Tay60b]. Temperature-Controlled [Fah64]. Temperatures [Bis60b]. their [Rai61b, VC66, Vat62]. Theoretical [Bar64, Sze69a]. Three [Far60a]. throughout [Ans68]. Tidal [Jon62b, VC62]. Tide [Bol62]. Tilapia [Gar61]. Time [NP64]. Toivo [Hal61]. Tomopteris [RS68]. Tool [dV69]. Torsion [Ree64]. Total [Hep66]. Tow [SCC68]. Towed [Bar64]. Towed-Net [Bar64]. trace [HMA66]. Traces [Rai64, VC62, VC66]. Transatlantic [Mat62]. Transport [Jør60]. Travin [Jon69, Sor64c]. Trawl [Lon60, Ric60a]. Trawlers [ZdV63]. Trawling [Bac66]. Trials [Bar68, CJG66, JM62]. Tropical [Hei69, Lon60, Lon64]. Trout [Kip62, Zar60, ZDG64]. trutta [Kip62]. Tuna [Mat62, SY60, Str61]. Tunas [Fin65]. Turbidity [LF69]. Two [Mat62, Wil69b, IJ62, eJPT66]. Types [Far60a, Rai61b]. typus [eJPT66].

Ultraviolet [Nie64]. Underwater [BeB60, Liv62, SY60]. Undisturbed [Cra65, Jør60]. Unfiltered [Jon63]. Untersuchungen [Sze68b]. Unusual [Rai61b]. upon [Jon62a, Sor64c]. Upper [Bis60b, Zar60]. Use [Gol68, HMA66, JM62, Kip62, Lov60, MS65, Pan62, RS67, SJI68, CNF68, Thr69]. Used [CHT61, Fin65, Ano65a]. Utilization [Men60].

V [Lar62, Mit63]. V. [eJPT66]. Valenciennes [Bai61, Bai63]. Value [Moy63, Ols64]. Values [Sit67]. Variations [Gul64c, TF65, VC66, Wil69b, dC66, Lin65]. Various [Bar68, Moy63]. Varying [Ril66b]. Veen [Las65]. Vellar [Bur69]. Venice [Vat61]. Venus [AL63, Ans64]. Vergleichende [Sze68b]. Vertical [BS69, BP65, MCM63, Sor60]. VI [Bow66]. Viability [Bow66]. VII [Ril66b]. Villefranche [Bra63]. Villefranche-sur-Mer [Bra63]. Vistula [Zar60]. Visual [Min68]. Volume [Ano60b, Ano61, Ano62, Ano64, Ano65b, Ano66d, Ano68d, Ano69b]. Volumes [Tra60]. vulgaris [EZ63, Hep66].

Waste [IS69]. Water [Bis60a, Bis62b, CW65, Day60, Fah64, Hol60b, Jon63, Jør60, MS68, RS67, Rum61, SP66, She63a, Sim64, Sit67, Sze69b, Tay60b, Vac62, Vat62, Dow64, BB68]. Waters [Gri60, TH68]. Weather [Jon69]. Weighing [Ree64]. Weight [HB67]. West [Lon60, Lon64, Lon65a, Roe69, eJPT66]. Western [Fur68, Hei69, Zij67]. Whale [Gul66b]. Whales [Roe69]. Whaling [Gul68b]. Whiting [GM64, Rou62]. whole [Wil67a]. Wild [Thr69]. William [Car69]. Wilson [Dow62]. Wind [Bol62, Cra60]. Winter [Aas62, BB68, Rou62]. Witch [Bow60]. Within [Pea62]. Work [Nie64].

XXIX [Ano65b]. XXV [Ano60b]. XXVI [Ano61]. XXVII [Ano62].

XXVIII [Ano64].

Year [IJ62, Ile67]. Year-Class [IJ62]. Yield [Gul68a, KQ68]. Young [Bis61a, Bis61b, Bis62b, Bis62a].

Zeiss [Thr69]. Zones [Bla69, Min68]. Zooplancton [Fur68]. Zooplankton [Jon62a, Lin65, MR61, RAL64, RKT67, Tra60, Fur68].

References

Aasen:1962:CBA

[Aas62] Olav Aasen. On the correlation between the arrival and spawning of the Norwegian winter herring. ICES Journal of Marine Science, 27(2):162–166, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/2/162/690296.

Ansell:1963:POR

[AL63] Alan D. Ansell and F. A. Loosmore. Preliminary observations on the relationship between growth, spawning and condition in experimental colonies of Venus mercenaria L. ICES Journal of Marine Science, 28(2):285–294, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/28/2/285/780270.

Albrechtsen:1968:DTO

[Alb68] K. Albrechtsen. A dyeing technique for otolith age reading. ICES Journal of Marine Science, 32(2):278–280, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
2/278/756682.

Albrechtsen:1969:NBF

[Alb69] K. Albrechtsen. A new bucket for filtration of microplankton. ICES Journal of Marine Science, 33(1):105–107, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
1/105/628728.

Allen:1967:SQM

[All67] K. Radway Allen. Some quick methods for estimating the effect on catch of changes in the size limit. ICES Journal of Marine Science, 31(1):111–126, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/1/111/843843.

Anonymous:1960:CBA

[Ano60a] Anonymous. “Current bibliography for aquatic sciences and fisheries”. ICES Journal of Marine Science, 25(2):231, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
2/231/623382.

Anonymous:1960:IVX

[Ano60b] Anonymous. Index to volume XXV. ICES Journal of Marine Science, 25(3):336–337, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/25/3/336/666350.

Anonymous:1961:IVX

[Ano61] Anonymous. Index to volume XXVI. ICES Journal of Marine Science, 26(3):357–358, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/26/3/357/872913.

Anonymous:1962:IVX

[Ano62] Anonymous. Index to volume XXVII. ICES Journal of Marine Science, 27(3):334–335, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/27/3/334/738965.

Anonymous:1964:IVX

[Ano64] Anonymous. Index to volume XXVIII. ICES Journal of Marine Science, 28(3):445–446, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/28/3/445/604011.

Anonymous:1965:GFM

[Ano65a] Anonymous. A guide to fish marks used by Members of the International Council for the Exploration of the Sea and by some Non-participant Countries. ICES Journal of Marine Science, 30 (1):87–160, December 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/1/87/642399.

Anonymous:1965:IVX

[Ano65b] Anonymous. Index to volume XXIX. ICES Journal of Marine Science, 29(3):354–355, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/29/3/354/689799.

Anonymous:1966:BR

[Ano66a] Anonymous. Books received. ICES Journal of Marine Science, 30(2):278, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/2/278/702902.

Anonymous:1966:GI

[Ano66b] Anonymous. Geographic index. ICES Journal of Marine Science, 30(3):436–438, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/3/436/690570.

Anonymous:1966:IJD

[Ano66c] Anonymous. Index of journal du conseil. ICES Journal of Marine Science, 30(3):361–362, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/30/3/361/690556.

Anonymous:1966:IV

[Ano66d] Anonymous. Index to volume 30. ICES Journal of Marine Science, 30(3):439–440, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/3/439/690573.

Anonymous:1966:RC

[Ano66e] Anonymous. References, citations. ICES Journal of Marine Science, 30(3):363–428, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/30/3/363/690559.

Anonymous:1966:SI

[Ano66f] Anonymous. Subject index. ICES Journal of Marine Science, 30 (3):429–432, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/3/429/690563.

Anonymous:1966:TI

[Ano66g] Anonymous. Taxonomic index. ICES Journal of Marine Science, 30(3):433–435, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/3/433/690566.

Anonymous:1967:BRa

[Ano67a] Anonymous. Books received. ICES Journal of Marine Science, 31(1):136, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/1/136/843874.

Anonymous:1967:BRb

[Ano67b] Anonymous. Books recevied. ICES Journal of Marine Science, 31(2):286, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/2/286/696760.

Anonymous:1968:BRa

[Ano68a] Anonymous. Books received. ICES Journal of Marine Science, 31(3):435, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/3/435/712503.

Anonymous:1968:BRb

[Ano68b] Anonymous. Books received. ICES Journal of Marine Science, 32(1):152, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/1/152/641400.

Anonymous:1968:BRc

[Ano68c] Anonymous. Books received. ICES Journal of Marine Science, 32(2):285, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/2/285/756701.

Anonymous:1968:IV

[Ano68d] Anonymous. Index to volume 31. ICES Journal of Marine Science, 31(3):436–437, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/3/436/712504.

Anonymous:1969:BR

[Ano69a] Anonymous. Books received. ICES Journal of Marine Science, 33(1):116, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/1/116/628737.

Anonymous:1969:IV

[Ano69b] Anonymous. Index to volume. ICES Journal of Marine Science, 32(3):444–445, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/3/444/714298.

Ansell:1964:SPG

[Ans64] Alan D. Ansell. Some parameters of growth of mature Venus mercenaria L. ICES Journal of Marine Science, 29(2):214–220, October 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
29/2/214/652943.

Ansell:1968:RGH

[Ans68] A. D. Ansell. The rate of growth of the hard clam (Mercenaria mercenaria L) throughout the geographical range. ICES Journal of Marine Science, 31(3):364–409, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/31/3/364/712492.

Armstrong:1962:NPP

[AW62] F. A. J. Armstrong and J. H. Wickstead. A note on the preservation of plankton samples with formalin. ICES Journal of Marine Science, 27(2):129–130, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/2/129/690279.

Backus:1966:PAD

[Bac66] Richard H. Backus. The “pinger” as an aid in deep trawling. ICES Journal of Marine Science, 30(2):270–277, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/30/
2/270/702899.

Bainbridge:1961:ELH

[Bai61] V. Bainbridge. The early life history of the Bonga, Ethmalosa dorsalis (Cuvier and Valenciennes). ICES Journal of Marine Science, 26(3):347–353, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/3/347/872889.

Bainbridge:1963:FFH

[Bai63] V. Bainbridge. The food, feeding habits and distribution of the Bonga Ethmalosa dorsalis (Cuvier & Valenciennes). ICES Journal of Marine Science, 28(2):270–284, September 1963. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/28/2/270/
780255.

Balls:1962:R

[Bal62] R. Balls. Review. ICES Journal of Marine Science, 27(3): 330–331, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/3/330/738952.

Barkley:1964:TET

[Bar64] Richard A. Barkley. The theoretical effectiveness of towed-net samplers as related to sampler size and to swimming speed of organisms. ICES Journal of Marine Science, 29(2):146–157, October 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/29/
2/146/652929.

Barker:1968:CTE

[Bar68] C. J. Barker. Consumption trials and edible fractions of various commercially important species of fish and shellfish. ICES Journal of Marine Science, 32(1):117–122, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/32/1/117/641389.

Burd:1965:SDH

[BB65] A. C. Burd and J. Bracken. Studies on the dunmore herring stock: 1. a population assessment. ICES Journal of Marine Science, 29(3):277–301, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/29/3/277/689771.

Banoub:1968:WDS

[BB68] M. W. Banoub and J. D. Burton. The winter distribution of silicate in Southampton Water. ICES Journal of Marine Science, 32 (2):201–208, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/2/201/756642.

Berthois:1960:MSM

[BeB60] L. Par Berthois and R. et Brenot. La morphologie sous-marine du talus du plateau continental entre le Sud de l’Irlande et le Cap Ortegal (Espagne): Mise au point des résultats obtenus. (French) [The underwater morphology of the Continental Shelf Slope between the South of Ireland and Cape Ortegal (Spain): Focus of the results obtained]. ICES Journal of Marine Science, 25(2):111–114, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
25/2/111/623317.

Beers:1964:AIP

[Bee64] John R. Beers. Ammonia and inorganic phosphorus excretion by the planktonic chaetognath, Sagitta hispida Conant. ICES Journal of Marine Science, 29(2):123–129, October 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/29/
2/123/652922.

Berg:1960:R

[Ber60] Kaj Berg. Reviews. ICES Journal of Marine Science, 25(2):230, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
25/2/230/623360.

Bernard:1965:LF

[Ber65] Francis Bernard. Louis Fage 1883–1964. ICES Journal of Marine Science, 29(3):233–236, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/29/3/233/689749.

Bertelsen:1968:HE

[Ber68] E. Bertelsen. Hermann Einarsson (1913–1966). ICES Journal of Marine Science, 32(1):3–6, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/32/1/3/641408.

Beyer:1967:SCM

[BFRT67] F. Beyer, E. Feyn, J. T. Ruud, and E. Totland. Stratified currents measured in the Oslofjord by means of a new, continuous depthcurrent recorder the bathyrheograph. ICES Journal of Marine Science, 31(1):5–26, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/1/5/843983.

Blaxter:1963:IES

[BH63] J. H. S. Blaxter and G. Hempel. The influence of egg size on herring larvae (Clupea harengus L.). ICES Journal of Marine Science, 28 (2):211–240, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/2/211/780198.

Bickmore:1967:R

[Bic67] D. P. Bickmore. Reviews. ICES Journal of Marine Science, 31(1):129–135, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/1/129/843862.

Birkett:1961:R

[Bir61] L. Birkett. Review. ICES Journal of Marine Science, 26(2):230– 231, April 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/26/2/230/651605.

Birkett:1969:CMS

[Bir69] L. Birkett. Calorimetry: Modification of a standard Bomb for small heat outputs. ICES Journal of Marine Science, 32(3):437–440, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
32/3/437/714295.

Bishai:1960:EWC

[Bis60a] H. M. Bishai. The effect of water currents on the survival and distribution of fish larvae. ICES Journal of Marine Science, 25 (2):134–146, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/2/134/623324.

Bishai:1960:ULT

[Bis60b] H. M. Bishai. Upper lethal temperatures for larval salmonids. ICES Journal of Marine Science, 25(2):129–133, March 1960.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
2/129/623322.

Bishai:1961:EPS

[Bis61a] H. M. Bishai. The effect of pressure on the survival and distribution of larval and young fish. ICES Journal of Marine Science, 26 (3):292–311, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/3/292/872867.

Bishai:1961:ESS

[Bis61b] H. M. Bishai. The effect of salinity on the survival and distribution of larval and young fish. ICES Journal of Marine Science, 26(2):166–179, April 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/2/166/651576.

Bishai:1962:RLYb

[Bis62a] H. M. Bishai. Reactions of larval and young salmonids to different hydrogen ion concentrations. ICES Journal of Marine Science, 27(2):181–191, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/2/181/690303.

Bishai:1962:RLYa

[Bis62b] H. M. Bishai. Reactions of larval and young salmonids to water of low oxygen concentration. ICES Journal of Marine Science, 27(2):167–180, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/2/167/690299.

Blacker:1969:CCZ

[Bla69] R. W. Blacker. Chemical composition of the zones in cod (Gadus morhua L.) otoliths. ICES Journal of Marine Science, 33(1): 107–108, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/1/107/628729.

Bolster:1962:IWT

[Bol62] G. C. Bolster. The influence of wind and tide on the catch by drift nets in the East Anglian herring fishery 1952–1956. ICES Journal of Marine Science, 27(2):150–161, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/27/2/150/690292.

Boltovskoy:1964:SOS

[Bol64]  Esteban Boltovskoy.  Seasonal occurrences of some living

foraminifera in Puerto Deseado (Patagonia, Argentina). ICES Journal of Marine Science, 29(2):136–145, October 1964. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/29/2/136/
652926.

Booker:1961:NST

[Boo61] P. G. Booker. New sea temperature measuring devices. ICES Journal of Marine Science, 26(2):133–147, April 1961. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/26/2/133/
651560.

Bougis:1963:CMA

[Bou63] Paul Bougis. Comparaison des méthods au diéthyldithiocarbamate de sodium (SDDC) et á la 2–2’-diquinolyle pour le dosage du cuivre en eau de mer. (French) [Comparison of sodium diethyldithiocarbamate (SDDC) and of 2–2’-diquinolyl methods for the determination of copper in seawater]. ICES Journal of Marine Science, 28 (2):171–174, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/2/171/780149.

Bowers:1960:GWG

[Bow60] A. B. Bowers. Growth of the witch (Glyptocephalus cynoglossus (L.)) in the Irish Sea. ICES Journal of Marine Science, 25 (2):168–176, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/2/168/623333.

Bowers:1966:MFC

[Bow66] A. B. Bowers. Marine fish culture in Britain: VI. The effect of the acclimatization of adult plaice to pond conditions on the viability of eggs and larvae. ICES Journal of Marine Science, 30(2):196–203, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
30/2/196/702869.

Blaxter:1965:ILS

[BP65] J. H. S. Blaxter and B. B. Parrish. The importance of light in shoaling, avoidance of nets and vertical migration by herring. ICES Journal of Marine Science, 30(1):40–57, December 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/30/1/40/642390.

Brock:1960:FSP

[BR60] Vernon E. Brock and Robert H. Riffenburgh. Fish schooling: a possible factor in reducing predation. ICES Journal of Marine Science, 25(3):307–317, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/25/3/307/666330.

Braconnot:1963:EDC

[Bra63] Jean Claude Braconnot. Étude du cycle annuel des salpes et dolioles en rade de Villefranche-sur-Mer. (French) [Study of the annual cycle of salps and dolioles in the Bay of Villefranche-sur-Mer]. ICES Journal of Marine Science, 28(1):21–36, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/28/1/21/611187.

Beers:1969:VDM

[BS69] J. R. Beers and G. L. Stewart. The vertical distribution of microzooplankton and some ecological observations. ICES Journal of Marine Science, 33(1):30–44, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/33/1/30/628765.

Beverton:1967:MPP

[BT67] R. J. H. Beverton and D. S. Tungate. A multi-purpose plankton sampler. ICES Journal of Marine Science, 31(2):145–157, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/31/
2/145/696717.

Burd:1963:SDF

[Bur63] A. C. Burd. On selection by the drifter fleets in the East Anglian herring fishery. ICES Journal of Marine Science, 28(1):91–120, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
28/1/91/611213.

Burton:1969:MVE

[Bur69] J. D. Burton. Molybdenum in the Vellar Estuary, Southern India. ICES Journal of Marine Science, 33(1):103–104, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
1/103/628727.

Coughlan:1964:DMD

[CA64] John Coughlan and Alan D. Ansell. A direct method for determining the pumping rate of siphonate bivalves. ICES Journal of Marine Science, 29(2):205–213, October 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/29/2/205/652941.

Candeias:1961:AMR

[Can61] Alberto Candeias. Alfredo de M. Ramalho 1894–1959. ICES Journal of Marine Science, 26(3):235–237, September 1961. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/26/3/235/
872835.

Carruthers:1969:WCH

[Car69] J. N. Carruthers. William Cuthbert Hodgson (1895–1968). ICES Journal of Marine Science, 33(1):7–10, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/33/1/7/628784.

Corey:1966:NQB

[CC66] Susan Corey and J. S. Craib. A new quantitative bottom sampler for meiof auna. ICES Journal of Marine Science, 30(3): 346–353, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/3/346/690548.

Chapman:1969:R

[Cha69] W. M. Chapman. Reviews. ICES Journal of Marine Science, 33 (1):114–116, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/1/114/628735.

Christensen:1964:BOT

[Chr64] Jergen Mailer Christensen. Burning of otoliths, a technique for age determination of soles and other fish. ICES Journal of Marine Science, 29(1):73–81, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/29/1/73/598992.

Colton:1961:ESM

[CHT61] John B. Colton, Jr., Kenneth A. Honey, and Robert F. Temple. The effectiveness of sampling methods used to study the distribution of larval herring in the Gulf of Maine. ICES Journal of Marine Science, 26(2):180–190, April 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/26/2/180/651581.

Cushing:1966:STM

[CJG66] D. H. Cushing, F. R. Harden Jones, and J. A. Gulland. Sea trials with modulation sector scanning sonar. ICES Journal of Marine Science, 30(3):324–342, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/30/3/324/690545.

Cushing:1968:UCC

[CNF68] D. H. Cushing, H. F. Nicholson, and G. P. Fox. The use of the Coulter counter for the determination of marine primary productivity. ICES Journal of Marine Science, 32(1):131–151, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
1/131/641397.

Cooper:1963:ER

[Coo63] L. H. N. Cooper. Essay review. ICES Journal of Marine Science, 28 (2):316–320, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/2/316/780308.

Cook:1968:R

[Coo68] J. C. Cook. Review. ICES Journal of Marine Science, 32(2): 283–284, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/2/283/756696.

Colebrook:1961:SCP

[CR61] J. M. Colebrook and G. A. Robinson. The seasonal cycle of the plankton in the North Sea and the North-Eastern Atlantic. ICES Journal of Marine Science, 26(2):156–165, April 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/26/
2/156/651569.

Craig:1960:NDC

[Cra60] R. E. Craig. A note on the dependence of catches on temperature and wind in the buchan pre-spawning herring fishery. ICES Journal of Marine Science, 25(2):185–190, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/25/2/185/623339.

Craib:1965:STS

[Cra65] J. S. Craib. A sampler for taking short undisturbed marine cores. ICES Journal of Marine Science, 30(1):34–39, December 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/30/
1/34/642387.

Culkin:1968:RAC

[Cul68] F. Culkin. Roland Arthur Cox. ICES Journal of Marine Science, 31(3):289–290, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/3/289/712475.

Cushing:1961:R

[Cus61] D. H. Cushing. Reviews. ICES Journal of Marine Science, 26 (3):355–356, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/3/355/872905.

Cushing:1962:AME

[Cus62a] D. H. Cushing. An alternative method of estimating the critical depth. ICES Journal of Marine Science, 27(2):131–140, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
2/131/690282.

Cushing:1962:R

[Cus62b] D. H. Cushing. Review. ICES Journal of Marine Science, 27 (3):331–333, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/3/331/738958.

Cushing:1967:R

[Cus67] D. H. Cushing. Reviews. ICES Journal of Marine Science, 31 (1):127–129, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/1/127/843855.

Cushing:1968:DSH

[Cus68a] D. H. Cushing. The Downs stock of herring during the period 1955– 1966. ICES Journal of Marine Science, 32(2):262–269, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
2/262/756673.

Cushing:1968:EAH

[Cus68b] D. H. Cushing. The East Anglian herring fishery in the eighteenth century. ICES Journal of Marine Science, 31(3):323–329, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/31/
3/323/712486.

Cushing:1968:GHC

[Cus68c] D. H. Cushing. Grazing by herbivorous copepodsin the sea. ICES Journal of Marine Science, 32(1):70–82, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/32/1/70/641413.

Cushing:1969:RSS

[Cus69] D. H. Cushing. The regularity of the spawning season of some fishes. ICES Journal of Marine Science, 33(1):81–92, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
1/81/628790.

Carlucci:1965:CBS

[CW65] A. F. Carlucci and P. M. Williams. Concentration of bacteria from sea water by bubble scavenging. ICES Journal of Marine Science, 30(1):28–33, December 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/1/28/642384.

Day:1960:BWT

[Day60] C. Godfrey Day. Bottom water temperature on browns ledge off Southern Massachusetts. ICES Journal of Marine Science, 25 (3):235–239, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/3/235/666314.

deCiechomski:1966:DLV

[dC66] Janina Dz. de Ciechomski. Development of the larvae and variations in the size of the eggs of the Argentine anchovy, Engraulis anchoita Hubbs and Marini. ICES Journal of Marine Science, 30 (3):281–290, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/3/281/690531.

deFigueiredo:1967:BNL

[dFT67] Maria José de Figueiredo and H. J. Thomas. On the biology of the Norway lobster, Nephrops norvegicus (L.). ICES Journal of Marine Science, 31(1):89–101, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/31/1/89/844000.

deGroot:1969:DSS

[dG69] S. J. de Groot. Digestive system and sensorial factors in relation to the feeding behaviour of flatfish (Pleuronectiformes). ICES Journal of Marine Science, 32(3):385–394, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/32/3/385/714275.

deMendiola:1969:FPA

[dM69] B. Rojas de Mendiola. The food of the Peruvian anchovy. ICES Journal of Marine Science, 32(3):433–434, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
3/433/714290.

Dow:1962:SSR

[Dow62] Robert L. Dow. Swimming speed of river herring Pomolobus pseudoharengus (Wilson). ICES Journal of Marine Science, 27 (1):77–80, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/1/77/581513.

Dow:1964:CAS

[Dow64] Robert L. Dow. A comparison among selected marine species of an association between sea water temperature and relative abundance. ICES Journal of Marine Science, 28(3):425–431, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/28/
3/425/604000.

Dutt:1969:SAG

[Dut69] S. Dutt. Studies on age and growth of the catfish Osteogneiosus militaris (Linn.). ICES Journal of Marine Science, 33(1): 109–110, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/1/109/628730.

deVeen:1967:PSS

[dV67] J. F. de Veen. On the phenomenon of soles (Solea solea L.) swimming at the surface. ICES Journal of Marine Science, 31 (2):207–236, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/2/207/696738.

deVeen:1969:APP

[dV69] J. F. de Veen. Abnormal pigmentation as a possible tool in the study of the populations of the plaice (Pleuronectes platessa L.). ICES Journal of Marine Science, 32(3):344–384, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
3/344/714273.

Troadec:1966:DAP

[eJPT66] Francis Poinsard et Jean-Paul Troadec. Détermination de l’Âge par la lecture des otolithes chez deux espèces de Sciaenidés ouestafricains: (Pseudotolithus senegalensis C. V. et Pseudotolithus typus Blkr.). (French) [Age determination by reading otoliths in two West African Sciaenidae species: (Pseudotolithus senegalensis CV and Pseudotolithus typus Blkr.)]. ICES Journal of Marine Science, 30(3):291–307, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/3/291/690538.

Engholm:1962:R

[Eng62] Basil Engholm. Review. ICES Journal of Marine Science, 27 (3):328–330, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/3/328/738946.

El-Zarka:1963:ASV

[EZ63] Salah-El-Din El-Zarka. Acclimatization of Solea vulgaris (Linn.) in Lake Quarun, Egypt. ICES Journal of Marine Science, 28 (1):126–136, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/1/126/611172.

Fahy:1964:TCS

[Fah64] William E. Fahy. A temperature-controlled salt-water circulating apparatus for developing fish eggs and larvae. ICES Journal of Marine Science, 28(3):364–384, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/28/3/364/603987.

Farris:1960:ETD

[Far60a] David A. Farris. The effect of three different types of growth curves on estimates of larval fish survival. ICES Journal of Marine Science, 25(3):294–306, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/25/3/294/666327.

Friethriksson:1960:NIF

[FAR60b] Árni Fridriksson, Olav Aasen, and Arne Revheim. New implements for fish tagging. ICES Journal of Marine Science, 25(2):158–167, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
25/2/158/623330.

Fink:1965:TEU

[Fin65] Bernard D. Fink. A technique, and the equipment used, for tagging tunas caught by the pole and line method. ICES Journal of Marine Science, 29(3):335–339, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/29/3/335/689782.

Fraser:1960:HO

[Fra60] J. H. Fraser. Helen Ogilvie 1880–1960. ICES Journal of Marine Science, 26(1):3–4, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/26/1/3/2182114.

Fraser:1964:R

[Fra64] J. H. Fraser. Review. ICES Journal of Marine Science, 29 (2):230, October 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/29/2/230/652949.

Furnestin:1964:PD

[Fur64] Furnestin. Pierre Desbrosses 1902–1962. ICES Journal of Marine Science, 28(3):325–326, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/28/3/325/603976.

Furnestin:1968:ZMB

[Fur68] Marie-Louise Furnestin. Le zooplancton de la Mediterranée (bassin occidental). Essai de synthèse. (French) [Zooplankton of the Mediterranean (Western Basin). Synthesis essay]. ICES Journal of Marine Science, 32(1):25–69, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/32/1/25/641404.

Garrod:1961:SCN

[Gar61] D. J. Garrod. The selection characteristics of nylon gill nets for Tilapia esculenta Graham. ICES Journal of Marine Science, 26(2):191–203, April 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/2/191/651583.

Grandperrin:1968:ALP

[GC68] R. Grandperrin and C. Caboche. Apercu sur l’action des procedes de conservation sur la biomasse d’organismes micronectoniques et macroplanctoniques. (French) [Overview of the action of conservation processes on the biomass of micronectonic and macroplankton organisms]. ICES Journal of Marine Science, 32 (2):209–215, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/2/209/756648.

German:1960:ISM

[Ger60] J. C. German. An inductive salinity meter for in situ measurements. ICES Journal of Marine Science, 26(1):21–32, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/26/
1/21/2182111.

Gulland:1961:SCM

[GH61] J. A. Gulland and D. Harding. The selection of Clarias mossambicus (Peters) by nylon gill nets. ICES Journal of Marine Science, 26(2):215–222, April 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/2/215/651593.

Gambell:1964:ADW

[GM64] R. Gambell and J. Messtorff. Age determination in the whiting (Merlangius merlangus L.) by means of the otoliths. ICES Journal of Marine Science, 28(3):393–404, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/28/3/393/603993.

Grandperrin:1969:ECF

[GM69] R. Grandperrin and A. Michel. Efficiences comparées de filets à plancton coniques de mêmes dimensions et de mailles différentes. I. Études générales. (French) [Comparative efficiencies of conical plankton nets of same dimensions and different meshes. I. General studies]. ICES Journal of Marine Science, 33(1):53–66, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
33/1/53/628777.

Goldman:1968:UAA

[Gol68] Charles R. Goldman. The use of absolute activity for eliminating serious errors in the measurement of primary productivity with C^{14}. ICES Journal of Marine Science, 32(2):172–179, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
2/172/756629.

Graham:1964:D

[Gra64] Michael Graham. A. T. A. Dobson 1885–1962. ICES Journal of Marine Science, 29(1):3–5, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/29/1/3/598976.

Gripenberg:1960:ABW

[Gri60] Stina Gripenberg. On the alkalinity of Baltic waters. ICES Journal of Marine Science, 26(1):5–20, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/26/1/5/2182121.

Griffiths:1968:EFS

[Gri68] P. G. Griffiths. An electronic fish scale proportioning system. ICES Journal of Marine Science, 32(2):280–282, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
2/280/756690.

Gulland:1961:EEC

[Gul61] J. A. Gulland. The estimation of the effect on catches of changes in gear selectivity. ICES Journal of Marine Science, 26(2):204–214, April 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
26/2/204/651590.

Gulland:1964:NIE

[Gul64a] J. A. Gulland. A note on the interim effects on catches of changes in gear selectivity. ICES Journal of Marine Science, 29(1):61–64, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
29/1/61/598984.

Gulland:1964:R

[Gul64b] J. A. Gulland. Review. ICES Journal of Marine Science, 29 (1):117–119, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/29/1/117/598970.

Gulland:1964:VSF

[Gul64c] J. A. Gulland. Variations in selection factors, and mesh differentials. ICES Journal of Marine Science, 29(2):158–165, October 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/29/
2/158/652932.

Gulland:1966:ATD

[Gul66a] J. A. Gulland. Analysis of target discrimination with sector scanning equipment. ICES Journal of Marine Science, 30(3): 343–345, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/3/343/690546.

Gulland:1966:ERA

[Gul66b] J. A. Gulland. The effect of regulation on Antarctic whale catches. ICES Journal of Marine Science, 30(3):308–315, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
30/3/308/690541.

Gulland:1968:CMY

[Gul68a] J. A. Gulland. The concept of the marginal yield from exploited fish stocks. ICES Journal of Marine Science, 32(2):256–261, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
32/2/256/756668.

Gulland:1968:MAW

[Gul68b] J. A. Gulland. The management of Antarctic whaling resources. ICES Journal of Marine Science, 31(3):330–341, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/31/
3/330/712488.

Gulland:1968:RCN

[Gul68c] J. A. Gulland. Recent changes in the North Sea plaice fishery. ICES Journal of Marine Science, 31(3):305, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/31/
3/305/712484.

Halim:1960:ONB

[Hal60] Y. Halim. Observations on the Nile Bloom of phytoplankton in the Mediterranean. ICES Journal of Marine Science, 26(1):57–67, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
26/1/57/2182122.

Halme:1961:THJ

[Hal61] Erkki Halme. Toivo Henrik Järvi 1877–1960. ICES Journal of Marine Science, 26(3):238–240, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/26/3/238/872836.

Hancock:1965:GEG

[Han65] D. A. Hancock. Graphical estimation of growth parameters. ICES Journal of Marine Science, 29(3):340–351, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/29/
3/340/689789.

Harris:1968:MMD

[Har68] J. G. K. Harris. A mathematical model describingthe possible behaviour of a copepodfeeding continuously in a relativelydense randomly distributed populationof algal cells. ICES Journal of Marine Science, 32(1):83–92, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/1/83/641416.

Haslett:1962:MDF

[Has62] R. W. G. Haslett. Measurement of the dimensions of fish to facilitate calculations of echo-strength in acoustic fish detection. ICES Journal of Marine Science, 27(3):261–269, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
3/261/738923.

Hempel:1961:EMM

[HB61] G. Hempel and J. H. S. Blaxter. The experimental modification of meristic characters in herring (Clupea harengus L.). ICES Journal of Marine Science, 26(3):336–346, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/26/
3/336/872883.

Hempel:1967:EWA

[HB67] G. Hempel and J. H. S. Blaxter. Egg weight in Atlantic herring (Clupea harengus L.). ICES Journal of Marine Science, 31 (2):170–195, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/2/170/696732.

Hancock:1967:EAG

[HE67] D. A. Hancock and E. Edwards. Estimation of annual growth in the edible crab (Cancer pagurus L.). ICES Journal of Marine Science, 31(2):246–264, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/2/246/696744.

Heinrich:1962:LHP

[Hei62] A. K. Heinrich. The life histories of plankton animals and seasonal cycles of plankton communities in the oceans. ICES Journal of Marine Science, 27(1):15–24, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/27/1/15/581491.

Heinrich:1969:TPC

[Hei69] A. K. Heinrich. On the tropical plankton communities in the Western Pacific. ICES Journal of Marine Science, 33(1):45–52, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
33/1/45/628775.

Hepper:1966:MCL

[Hep66] B. T. Hepper. Measurement of carapace length and total length in the crawfish, Palinurus vulgaris. ICES Journal of Marine Science, 30(3):316–323, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/3/316/690544.

Harvey:1966:NDS

[HG66] J. G. Harvey and W. J. Gould. A note on the design of sea-surface drifters. ICES Journal of Marine Science, 30(3):358–360, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
30/3/358/690554.

Holm-Hansen:1965:FDC

[HHLHS65] Osmund Holm-Hansen, Carl J. Lorenzen, Robert W. Holmes, and John D. H. Strickland. Fluorometric determination of chlorophyll. ICES Journal of Marine Science, 30(1):3–15, December 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/30/
1/3/642386.

Hiemstra:1962:CTA

[Hie62] Wim H. Hiemstra. A correlation table as an aid for identifying pelagic fish eggs in plankton samples. ICES Journal of Marine Science, 27(1):100–108, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/27/1/100/581477.

Holden:1964:FSS

[HM64] M. J. Holden and P. S. Meadows. The fecundity of the spurdog (Squalus acanthias L.). ICES Journal of Marine Science, 28(3):418–424, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/3/418/603999.

Hunter:1968:FEA

[HM68] John R. Hunter and Charles T. Mitchell. Field experiments on the attraction of pelagic fish to floating objects. ICES Journal of Marine Science, 31(3):427–434, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/31/3/427/712499.

Haven:1966:UFP

[HMA66] Dexter S. Haven and R. Morales-Alamo. Use of fluorescent particles to trace oyster biodeposits in marine sediments. ICES Journal of Marine Science, 30(2):267–269, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/30/2/267/702889.

Hoglund:1969:ARM

[Hög69a] Hans Höglund. Arvid R. Molander (1886–1965). ICES Journal of Marine Science, 33(1):11–13, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/33/1/11/628731.

Hoglund:1969:KAA

[Hög69b] Hans Höglund. Karl Andreas Andersson “K. A.” (1875–1968). ICES Journal of Marine Science, 33(1):3–6, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
1/3/628760.

Holden:1960:ECG

[Hol60a] M. J. Holden. Evidence of cod (Gadus morhua L.) migrations from the Norway Coast to the Faroese Islands. ICES Journal of Marine Science, 26(1):68–72, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/26/1/68/2182125.

Holt:1960:WTC

[Hol60b] S. J. Holt. Water temperature and cod growth-rate. ICES Journal of Marine Science, 25(2):225–228, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/25/2/225/623353.

Holden:1966:FSS

[Hol66] M. J. Holden. The food of the spurdog, Squalus acanthias (L.). ICES Journal of Marine Science, 30(2):255–266, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/30/
2/255/702884.

Hopkins:1968:CNC

[Hop68] Thomas L. Hopkins. Carbon and nitrogen content of fresh and preserved Nematoscelis dijficilis, a euphausiid crustacean. ICES Journal of Marine Science, 31(3):300–304, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/31/
3/300/712481.

Hisard:1969:DDN

[HP69] Philippe Hisard and Bernard Piton. La distribution du nitrite dans le système des courants équatoriaux de l’ocean Pacifique, à 170^{o}E. (French) [Distribution of nitrite in the Equatorial Current system of the Pacific Ocean, at 170 ^{o}E]. ICES Journal of Marine Science, 32(3):303–317, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/3/303/714265.

Hulburt:1960:PSS

[HRG60] E. M. Hulburt, J. H. Ryther, and R. R. L. Guillard. The phytoplankton of the Sargasso Sea off Bermuda. ICES Journal of Marine Science, 25(2):115–128, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/25/2/115/623319.

Iles:1962:CTA

[IJ62] T. D. Iles and P. O. Johnson. The correlation table analysis of a sprat (Clupea sprattus L.) year-class to separate two groups differing in growth characteristics. ICES Journal of Marine Science, 27 (3):287–303, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/3/287/738931.

Iles:1964:DMS

[Ile64] T. D. Iles. The duration of maturation stages in herring. ICES Journal of Marine Science, 29(2):166–188, October 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/29/
2/166/652934.

Iles:1967:GSN

[Ile67] T. D. Iles. Growth studies on North Sea herring: I. The second year’s growth (I-group) of East Anglian herring, 1939–63. ICES Journal of Marine Science, 31(1):56–76, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/31/1/56/843989.

Iles:1968:GSN

[Ile68] T. D. Iles. Growth studies on North Sea herring: II. 0-Group growth of East Anglian herring. ICES Journal of Marine Science, 32(1):98–116, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/1/98/641422.

Isaacs:1969:SMP

[IS69] John D. Isaacs and Walter R. Schmitt. Stimulation of marine productivity with waste heat and mechanical power. ICES Journal of Marine Science, 33(1):20–29, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/33/1/20/628750.

J:1962:CRR

[J.62] F. R. H. J. Cooperative research report. ICES Journal of Marine Science, 27(2):124, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/2/124/690274.

Jones:1963:DPE

[JH63] P. G. W. Jones and S. M. Haq. The distribution of Phaeocystis in the Eastern Irish Sea. ICES Journal of Marine Science, 28(1):8–20, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
28/1/8/611207.

Jones:1962:UES

[JM62] F. R. Harden Jones and B. S. McCartney. The use of electronic sector-scanning sonar for following the movements of fish shoals: Sea trials on R.R.S. “Discovery II”. ICES Journal of Marine Science, 27(2):141–149, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/2/141/690286.

Jones:1960:MSA

[Jon60] Rodney Jones. Mesh selection and apparent growth of haddock. ICES Journal of Marine Science, 25(2):177–184, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
2/177/623337.

Jones:1962:EIE

[Jon62a] Everet C. Jones. Evidence of an Island effect upon the standing crop of zooplankton near the Marquesas Islands, Central Pacific. ICES Journal of Marine Science, 27(3):223–231, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
3/223/738905.

Jones:1962:FOM

[Jon62b] F. R. Harden Jones. Further observations on the movements of herring (Clupea harengus L.) shoals in relation to the tidal current. ICES Journal of Marine Science, 27(1):52–76, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
1/52/581509.

Jones:1963:ECS

[Jon63] P. G. W. Jones. The effect of chloroform on the soluble inorganic phosphate content of unfiltered sea water. ICES Journal of Marine Science, 28(1):3–7, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/1/3/611192.

Jones:1969:SCP

[Jon69] D. H. Jones. Some characteristics of the pelagic redfish (Sebastes mentella Travin) from Weather Station Alfax. ICES Journal of Marine Science, 32(3):395–412, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/32/3/395/714279.

Jorgensen:1960:EPR

[Jør60] C. Barker Jørgensen. Efficiency of particle retention and rate of water transport in undisturbed lamellibranchs. ICES Journal of Marine Science, 26(1):94–116, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/26/1/94/2182132.

Kabata:1963:ICS

[Kab63] Z. Kabata. Incidence of coccidioses in Scottish herring (Clupea harengus L.). ICES Journal of Marine Science, 28(2):201–210, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/28/2/201/780197.

Kipling:1962:USB

[Kip62] Charlotte Kipling. The use of the scales of the brown trout (Salmo trutta L.) for the back-calculation of growth. ICES Journal of Marine Science, 27(3):304–315, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/27/3/304/738937.

Korringa:1967:RS

[Kor67] P. Korringa. Ragnar Spärck (1896–1965). ICES Journal of Marine Science, 31(1):3–4, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/1/3/843960.

Korringa:1969:R

[Kor69] P. Korringa. Reviews. ICES Journal of Marine Science, 32 (3):441–442, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/3/441/714296.

Kutty:1968:EOA

[KQ68] M. Krishnan Kutty and S. Z. Qasim. The estimation of optimum age of exploitation and potential yield in fish populations. ICES Journal of Marine Science, 32(2):249–255, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
32/2/249/756663.

Krey:1961:DIM

[Kre61] Johannes Krey. Der detritus im meere. ICES Journal of Marine Science, 26(3):263–280, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/26/3/263/872855.

Larsen:1962:CVO

[Lar62] Knud Larsen. C. V. Otterstrøm 1881–1962. ICES Journal of Marine Science, 27(3):219–220, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/27/3/219/738899.

Lassig:1965:IVV

[Las65] J. Lassig. An improvement to the van Veen bottom grab. ICES Journal of Marine Science, 29(3):352–353, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/29/
3/352/689794.

Lumby:1965:SCO

[LE65] J. R. Lumby and D. J. Ellett. Some considerations on oceanographic observations. ICES Journal of Marine Science, 29(3):237– 248, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/29/3/237/689750.

Lee:1969:FAT

[LF69] Arthur J. Lee and A. R. Folkard. Factors affecting turbidity in the Southern North Sea. ICES Journal of Marine Science, 32(3):291–302, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/3/291/714252.

Lindroth:1962:GA

[Lin62] Arne Lindroth. Gunnar Alm 1889–1962. ICES Journal of Marine Science, 27(3):221–222, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/27/3/221/738901.

Lindroth:1963:BSR

[Lin63] Arne Lindroth. The Body/ scale relationship in Atlantic salmon (Salmo salar L.) a preliminary report. ICES Journal of Marine Science, 28(1):137–152, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/28/1/137/611177.

Linford:1965:BSM

[Lin65] Eileen Linford. Biochemical studies on marine zooplankton: II. Variations in the lipid content of some Mysidacea. ICES Journal of Marine Science, 30(1):16–27, December 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/30/1/16/642383.

Lisitzin:1966:GG

[Lis66] Eugenie Lisitzin. Gunnar Granqvist 1888–1965. ICES Journal of Marine Science, 30(3):279–280, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/30/3/279/690529.

Livingstone:1962:UTO

[Liv62] Robert Livingstone, Jr. Underwater television observations of haddock (Melanogrammus aeglefinus [Linnaeus]) in the cod-end. ICES Journal of Marine Science, 27(1):43–48, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/27/1/43/581505.

Longhurst:1960:MSF

[Lon60] Alan R. Longhurst. Mesh selection factors in the trawl fishery off Tropical West Africa. ICES Journal of Marine Science, 25 (3):318–325, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/3/318/666335.

Longhurst:1964:BST

[Lon64] Alan R. Longhurst. Bionomics of the Sciaenidae of Tropical West Africa. ICES Journal of Marine Science, 29(1):93–114, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/29/
1/93/598997.

Longhurst:1965:BWA

[Lon65a] Alan R. Longhurst. The biology of West African polynemid fishes. ICES Journal of Marine Science, 30(1):58–74, December 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/30/
1/58/642394.

Longhurst:1965:SFR

[Lon65b] Alan R. Longhurst. A survey of the fish resources of the Eastern Gulf of Guinea. ICES Journal of Marine Science, 29(3):302–334, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
29/3/302/689776.

Lovegrove:1960:IFS

[Lov60] T. Lovegrove. An improved form of sedimentation apparatus for use with an inverted microscope. ICES Journal of Marine Science, 25(3):279–284, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/3/279/666319.

Love:1962:MCN

[Lov62] R. M. Love. The measurement of ‘condition’ in North Sea cod. ICES Journal of Marine Science, 27(1):34–42, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
1/34/581502.

Lasker:1962:OCO

[LT62] Reuben Lasker and Gail H. Theilacker. Oxygen consumption and osmoregulation by single Pacific sardine eggs and larvae (Sardinops caerulea Girard). ICES Journal of Marine Science, 27(1):25–33, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/1/25/581496.

Macer:1967:NBP

[Mac67] C. T. Macer. A new bottom-plankton sampler. ICES Journal of Marine Science, 31(2):158–163, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/31/2/158/696722.

Marak:1960:FHL

[Mar60] Robert R. Marak. Food habits of larval cod, haddock, and coalfish in the Gulf of Maine and Georges Bank area. ICES Journal of Marine Science, 25(2):147–157, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/25/2/147/623328.

Marr:1963:SNP

[Mar63] James Marr. Some notes on the preservation of marine animals. ICES Journal of Marine Science, 28(1):121–125, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/28/
1/121/611169.

Mather:1962:TMT

[Mat62] Frank J. Mather. Transatlantic migration of two large bluefin tuna. ICES Journal of Marine Science, 27(3):325–327, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
27/3/325/738942.

Miller:1963:SVD

[MCM63] David Miller, John B. Colton, Jr., and Robert R. Marak. A study of the vertical distribution of larval hadddock. ICES Journal of Marine Science, 28(1):37–49, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/28/1/37/611194.

Menzel:1960:UFB

[Men60] David W. Menzel. Utilization of food by a Bermuda reef fish, Epinephelus guttatus. ICES Journal of Marine Science, 25(2):216– 222, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/25/2/216/623345.

Millar:1964:BGC

[Mil64] R. H. Millar. Breeding and gonadial cycle of oysters in Loch Ryan, Scotland. ICES Journal of Marine Science, 28(3):432–439, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/28/
3/432/604003.

Mina:1968:NPV

[Min68] M. V. Mina. A note on a problem in the visual qualitative evaluation of otolith zones. ICES Journal of Marine Science, 32(1):93–97, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
32/1/93/641418.

Mitson:1963:MFC

[Mit63] R. B. Mitson. Marine fish culture in Britain: V. An electronic device for counting the nauplii of Artemia salina L. ICES Journal of Marine Science, 28(2):262–269, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/28/
2/262/780236.

Mahnken:1967:FEH

[MJ67] C. V. W. Mahnken and J. W. Jossi. Flume experiments on the hydrodynamics of plankton nets. ICES Journal of Marine Science, 31(1):38–45, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/1/38/843973.

Morcos:1968:SSS

[Mor68] Selim A. Morcos. Substandard seawater of any salinity for chlorinity determination. ICES Journal of Marine Science, 31(3): 291–299, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/3/291/712478.

Moyse:1963:CVV

[Moy63] John Moyse. A comparison of the value of various flagellates and diatoms as food for barnacle larvae. ICES Journal of Marine Science, 28(2):175–187, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/2/175/780162.

Muzinic:1960:SOB

[MP60] R. Muzinic and B. B. Parrish. Some observations on the body proportions of North Sea autumn spawning herring. ICES Journal of Marine Science, 25(2):191–203, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/25/2/191/623341.

McNeely:1961:SSD

[MP61] Richard L. McNeely and Walter T. Pereyra. A simple screening device for the separation of benthic samples at sea. ICES Journal of Marine Science, 26(3):259–262, September 1961. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/26/3/259/
872851.

McAllister:1960:PPF

[MPS60] C. D. McAllister, T. R. Parsons, and J. D. H. Strickland. Primary productivity and fertility at Station “P” in the North-East Pacific Ocean. ICES Journal of Marine Science, 25(3):240–259, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
3/240/666315.

Menzel:1961:ZSS

[MR61] David W. Menzel and John H. Ryther. Zooplankton in the Sargasso Sea off Bermuda and its relation to organic production. ICES Journal of Marine Science, 26(3):250–258, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
26/3/250/872846.

Millar:1965:UDC

[MS65] R. H. Millar and J. M. Scott. The use of a direct counting technique for bacteria in marine culture experiments. ICES Journal of Marine Science, 29(3):253–255, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/29/3/253/689759.

Millar:1968:EWQ

[MS68] R. H. Millar and J. M. Scott. An effect of water quality on the growth of cultured larvae of the oyster Ostrea edulis L. ICES Journal of Marine Science, 32(1):123–130, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/32/1/123/641392.

Mauchline:1966:SCB

[MT66] J. Mauchline and W. L. Templeton. Strontium, calcium and barium in marine organisms from the Irish Sea. ICES Journal of Marine Science, 30(2):161–170, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/30/2/161/702857.

Mitson:1961:AMC

[MW61] R. B. Mitson and R. J. Wood. An automatic method of counting fish echoes. ICES Journal of Marine Science, 26(3):281–291, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/26/3/281/872862.

Nielsen:1960:R

[Nie60a] E. Steemann Nielsen. Reviews. ICES Journal of Marine Science, 25(3):332, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/3/332/666343.

Nielsen:1960:PRT

[Nie60b] Jørgen Nielsen. Preliminary results of tagging experiments with herring (Clupea harengus L.) in Greenland. ICES Journal of Marine Science, 26(1):73–79, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/26/1/73/2182126.

Nielsen:1964:CMP

[Nie64] E. Steemann Nielsen. On a complication in marine productivity work due to the influence of ultraviolet light. ICES Journal of Marine Science, 29(2):130–135, October 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/29/2/130/652924.

Nielsen:1964:TCA

[NP64] E. Steemann Nielsen and Tai Soo Park. On the time course in adapting to low light intensities in marine phytoplankton. ICES Journal of Marine Science, 29(1):19–24, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/29/1/19/598973.

Olafsson:1967:AF

[Ola67] David Olafsson. Arni Frioriksson (1898–1966). ICES Journal of Marine Science, 31(2):139–144, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/31/2/139/696715.

Olson:1964:SVF

[Ols64] F. C. W. Olson. The survival value of fish schooling. ICES Journal of Marine Science, 29(1):115–116, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/29/1/115/598968.

Pruter:1962:ADG

[PA62] Alonzo T. Pruter and Dayton L. Alverson. Abundance, distribution, and growth of flounders in the South-Eastern Chukchi Sea. ICES Journal of Marine Science, 27(1):81–99, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
1/81/581514.

Paasche:1960:RBP

[Paa60] E. Paasche. On the relationship between primary production and standing stock of phytoplankton. ICES Journal of Marine Science, 26(1):33–48, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/1/33/2182115.

Pantulu:1962:UPS

[Pan62] V. Rayappa Pantulu. On the use of pectoral spines for the determination of age and growth of Pangasius pangasius (Hamilton Buch). ICES Journal of Marine Science, 27(2):192–216, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
2/192/690307.

Pantulu:1963:SAG

[Pan63] V. Rayappa Pantulu. Studies on the age and growth, fecundity and spawning of Osteogeneiosus militaris (Linn.). ICES Journal of Marine Science, 28(2):295–315, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/28/2/295/780290.

Pantulu:1969:AGR

[Pan69] V. R. Pantulu. Age and growth: a reply. ICES Journal of Marine Science, 33(1):111–113, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/33/1/111/628732.

Patten:1964:RDP

[Pat64a] Bernard C. Patten. The rational decision process in salmon migration. ICES Journal of Marine Science, 28(3):410–417, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/28/
3/410/603997.

Patten:1964:SM

[Pat64b] Bernard C. Patten. Salmon migratrion. ICES Journal of Marine Science, 28(3):440–444, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/28/3/440/604007.

Pearcy:1962:DOD

[Pea62] William G. Pearcy. Distribution and origin of demersal eggs within the order Pleuronectiformes. ICES Journal of Marine Science, 27 (3):232–235, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/3/232/738909.

Patten:1964:SEC

[PNYR64] Bernard C. Patten, John J. Norcross, David K. Young, and Charles L. Rutherford. Some experimental characteristics of dark and light bottles. ICES Journal of Marine Science, 28(3):335–353, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
28/3/335/603979.

Preston:1963:R

[Pre63] A. Preston. Review. ICES Journal of Marine Science, 28 (2):321, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/2/321/780312.

Pope:1967:SBO

[PT67] J. A. Pope and H. J. Thomas. Some biometric observations on Nephrops norvegicus (L.). ICES Journal of Marine Science, 31 (2):265–271, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/2/265/696748.

Postuma:1965:IHN

[PZD65] K. H. Postuma, J. J. Zijlstra, and N. Das. On the immature herring of the North Sea. ICES Journal of Marine Science, 29(3):256–276, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
29/3/256/689766.

Raitt:1961:OSS

[Rai61a] D. F. S. Raitt. Otolith studies of Southern North Sea herring. ICES Journal of Marine Science, 26(3):312–328, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
26/3/312/872873.

Raitt:1961:SUO

[Rai61b] D. F. S. Raitt. Some unusual otolith types on the Bløden Ground in 1958 and their subsequent occurrence in the adult fisheries. ICES Journal of Marine Science, 26(3):329–335, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/26/
3/329/872879.

Raitt:1962:NPC

[Rai62] D. F. S. Raitt. A note on the pyloric caecae of Gadus poutassou Risso. ICES Journal of Marine Science, 27(1):49–51, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
1/49/581508.

Raitt:1964:SRL

[Rai64] D. F. S. Raitt. Scottish redfish larval investigations in 1962 with some observations on mid-oceanic echo-traces. ICES Journal of Marine Science, 29(1):65–72, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/29/1/65/598989.

Raymont:1964:BSM

[RAL64] J. E. G. Raymont, J. Austin, and Eileen Linford. Biochemical studies on marine zooplankton: I. The biochemical composition of Neomysis integer. ICES Journal of Marine Science, 28(3):354–363, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
28/3/354/603983.

Reeve:1964:STM

[Ree64] M. R. Reeve. A simple torsion microbalance for weighing small animals. ICES Journal of Marine Science, 28(3):385–392, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/28/
3/385/603990.

Raitt:1967:FRS

[RH67] D. F. S. Raitt and W. B. Hall. On the fecundity of the redfish, Sebastes marinus (L.). ICES Journal of Marine Science, 31 (2):237–245, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/2/237/696742.

Richardson:1960:OSN

[Ric60a] I. D. Richardson. Observations on the size and numbers of herring taken by the herring trawl. ICES Journal of Marine Science, 25(2):204–215, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/2/204/623344.

Ricker:1960:R

[Ric60b] W. E. Ricker. Review. ICES Journal of Marine Science, 26(1): 125–129, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/1/125/2182106.

Rice:1963:FIS

[Ric63] A. L. Rice. The food of the Irish Sea herring in 1961 and 1962. ICES Journal of Marine Science, 28(2):188–200, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
28/2/188/780190.

Riffenburgh:1960:NME

[Rif60] Robert H. Riffenburgh. A new method for estimating parameters for the Gompertz growth curve. ICES Journal of Marine Science, 25(3):285–293, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/3/285/666324.

Riley:1966:LLM

[Ril66a] John D. Riley. Liquid latex marking technique for small fish. ICES Journal of Marine Science, 30(3):354–357, November 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/30/
3/354/690551.

Riley:1966:MFC

[Ril66b] John D. Riley. Marine fish culture in Britain: VII. Plaice (Pleuronectes platessa L.) post-larval feeding on Artemia salina L. nauplii and the effects of varying feeding levels. ICES Journal of Marine Science, 30(2):204–221, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/30/2/204/702873.

Raymont:1967:BSM

[RKT67] J. E. G. Raymont, S. Krishnaswamy, and J. Tundisi. Biochemical studies on marine zooplankton: IV. Investigations on succinic dehydrogenase activity in zooplankton with special reference to Neomysis integer. ICES Journal of Marine Science, 31(2):164–169, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
31/2/164/696726.

Roe:1969:FFH

[Roe69] H. S. J. Roe. The food and feeding habits of the sperm whales (Physeter catodon L.) taken off the West Coast of Iceland. ICES Journal of Marine Science, 33(1):93–102, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/33/1/93/628792.

Rosseland:1962:SS

[Ros62] Leiv Rosseland. Sven Sømme 1904–1961. ICES Journal of Marine Science, 27(2):119–120, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/27/2/119/690268.

Rothschild:1967:CGM

[Rot67] Brian J. Rothschild. Competition for gear in a multiple-species fishery. ICES Journal of Marine Science, 31(1):102–110, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/31/
1/102/843833.

Rout:1962:SOW

[Rou62] D. W. R. Rout. Some observations on the whiting (Gadus merlangus L.) of the inshore winter fishery off Lowestoft. ICES Journal of Marine Science, 27(3):316–324, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/27/3/316/738939.

Rakusa-Suszczewski:1967:UCC

[RS67] S. Rakusa-Suszczewski. The use of chaetognath and copepod population age-structures as an indication of similarity between water masses. ICES Journal of Marine Science, 31(1):46–55, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/31/
1/46/843978.

Rakusa-Suszczewski:1968:PCT

[RS68] S. Rakusa-Suszczewski. Predation of Chaetognatha by Tomopteris helgolandica Greff. ICES Journal of Marine Science, 32(2): 226–231, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/2/226/756656.

Riley:1963:MFC

[RT63] John D. Riley and Graham T. Thacker. Marine fish culture in Britain III. Plaice (Pleuronectes platessa (L)) rearing in closed circulation at Lowestoft, 1961. ICES Journal of Marine Science, 28(1):80–90, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/1/80/611210.

Rumney:1961:SWT

[Rum61] George R. Rumney. A shallow water temperature profile recording system. ICES Journal of Marine Science, 26(3):243–249, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
26/3/243/872843.

Ruud:1961:AND

[Ruu61] Johan T. Ruud. Alf Nicolai Dannevig 1886–1960. ICES Journal of Marine Science, 26(3):241–242, September 1961. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/26/3/241/
872840.

Roger:1968:TDG

[RW68] C. Roger and B. Wauthy. Sur une technique de détermination de groupes de tailles, applicable a l’Étude de certains organisms planctoniques. (French) [On a size group determination technique, applicable to the study of certain planktonic organisms]. ICES Journal of Marine Science, 32(2):216–225, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
2/216/756652.

Ryland:1966:ODL

[Ryl66] J. S. Ryland. Observations on the development of larvae of the plaice, Pleuronectes platessa L., in aquaria. ICES Journal of Marine Science, 30(2):177–195, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/30/2/177/702866.

Saila:1969:R

[Sai69] S. B. Saila. Reviews. ICES Journal of Marine Science, 32 (3):442–443, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/3/442/714297.

Sanders:1969:MDE

[San69] M. J. Sanders. A method of directly estimating natural mortality and initial tagging mortality applicable to certain exploited mollusc populations. ICES Journal of Marine Science, 32(3):416–418, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
32/3/416/714284.

Smith:1968:CFE

[SCC68] Paul E. Smith, Robert C. Counts, and Robert I. Clutter. Changes in filtering efficiency of plankton nets due to clogging under tow. ICES Journal of Marine Science, 32(2):232–248, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
2/232/756661.

Saila:1968:MBB

[SF68] S. B. Saila and J. M. Flowers. Movements and behaviour of berried female lobsters displaced from offshore areas to Narragansett Bay, Rhode Island. ICES Journal of Marine Science, 31 (3):342–351, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/31/3/342/712489.

Shelbourne:1963:MFCb

[She63a] J. E. Shelbourne. Marine fish culture in Britain II. A plaice rearing experiment at Port Erin, Isle of Man, during 1960, in open sea water circulation. ICES Journal of Marine Science, 28(1):70–79, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
28/1/70/611202.

Shelbourne:1963:MFCc

[She63b] J. E. Shelbourne. Marine fish culture in Britain: IV. High survivals of metamorphosed plaice during salinity experiments in open circulation at Port Erin, Isle of Man, 1961. ICES Journal of Marine Science, 28(2):246–261, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/28/2/246/780223.

Sheldon:1968:EHP

[She68] R. W. Sheldon. The effect of high population density on the growth and mortality of oysters (Ostrea edulis). ICES Journal of Marine Science, 31(3):352–363, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/31/3/352/712491.

Simkiss:1964:IES

[Sim64] K. Simkiss. The inhibitory effects of some metabolites on the precipitation of calcium carbonate from artificial and natural sea water. ICES Journal of Marine Science, 29(1):6–18, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/29/
1/6/598982.

Sitaramaiah:1967:WNC

[Sit67] P. Sitaramaiah. Water, nitrogen and calorific values of freshwater organisms. ICES Journal of Marine Science, 31(1):27–30, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/31/
1/27/843881.

Strasburg:1968:USS

[SJI68] Donald W. Strasburg, Everet C. Jones, and Robert T. B. Iversen. Use of a small submarine for biological and oceanographic research. ICES Journal of Marine Science, 31(3):410–426, January 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/31/
3/410/712494.

Stubbs:1962:AAS

[SL62] A. R. Stubbs and R. G. G. Lawrie. Asdic as an aid to spawning ground investigations. ICES Journal of Marine Science, 27 (3):248–260, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/3/248/738919.

Sundnes:1965:BDE

[SLI65] G. Sundnes, H. Leivestad, and O. Iversen. Buoyancy determination of eggs from the cod (Gadus morhua L.). ICES Journal of Marine Science, 29(3):249–252, April 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/29/3/249/689757.

Smed:1961:R

[Sme61] Jens Smed. Reviews. ICES Journal of Marine Science, 26(3): 354–355, September 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/3/354/872895.

Smed:1963:IOD

[Sme63] Jens Smed. ICES Oceanographic data lists. ICES Journal of Marine Science, 28(1):167, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/28/1/167/611184.

Smed:1968:SHI

[Sme68] Jens Smed. The service hydrographique of the International Council for The Exploration of the Sea. ICES Journal of Marine Science, 32(2):155–171, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/2/155/756624.

Smith:1968:OAR

[Smi68] S. Wiedemann Smith. Otolith age reading by means of surface structure examination. ICES Journal of Marine Science, 32(2): 270–277, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/2/270/756678.

Sorokin:1960:VDP

[Sor60] J. I. Sorokin. Vertical distribution of phytoplankton and the primary organic production in the sea. ICES Journal of Marine Science, 26(1):49–56, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/26/1/49/2182119.

Sorokin:1964:PPB

[Sor64a] J. I. Sorokin. On the primary production and bacterial activities in the Black Sea. ICES Journal of Marine Science, 29(1):41–60, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
29/1/41/598979.

Sorokin:1964:QSM

[Sor64b] J. I. Sorokin. A quantitative study of the microflora in the Central Pacific Ocean. ICES Journal of Marine Science, 29(1):25–40, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
29/1/25/598975.

Sorokin:1964:IFU

[Sor64c] V. P. Sorokin. On the influence of the fishery upon the population structure of redfish (Sebastes marinus L. and Sebastes mentella Travin). ICES Journal of Marine Science, 28(3):405–409, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/28/
3/405/603994.

Schumann:1966:JCS

[SP66] George O. Schumann and Herbert C. Perkins. A jar for culturing small pelagic marine organisms in running water during shipboard or laboratory studies. ICES Journal of Marine Science, 30(2):171–176, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/2/171/702861.

Sparck:1962:AFB

[Spä62] R. Spärck. Anton Frederik Bruun. 14. December 1901–13. December 1961. ICES Journal of Marine Science, 27(2):121–123, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
2/121/690271.

Spencer:1964:EPP

[Spe64] C. P. Spencer. The estimation of phytoplankton pigments. ICES Journal of Marine Science, 28(3):327–334, March 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/28/
3/327/603977.

Shelbourne:1963:MFCa

[SRT63] J. E. Shelbourne, J. D. Riley, and G. T. Thacker. Marine fish culture in Britain I. Plaice rearing in closed circulation at Lowestoft, 1975–1960. ICES Journal of Marine Science, 28(1):50–69, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
28/1/50/611199.

Saila:1963:RMO

[SS63] Saul B. Saila and Raymond A. Shappy. Random movement and orientation in salmono migration. ICES Journal of Marine Science, 28(1):153–166, June 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/1/153/611180.

Stefansson:1968:NPR

[Ste68] U. Stefánsson. Nitrate-phosphate relationships in the Irminger Sea. ICES Journal of Marine Science, 32(2):188–200, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
2/188/756637.

Stride:1960:R

[Str60a] A. H. Stride. Reviews. ICES Journal of Marine Science, 25(2):229, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
25/2/229/623356.

Strubberg:1960:R

[Str60b] E. Strubberg. Reviews. ICES Journal of Marine Science, 25 (3):332–335, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/3/332/666347.

Strasburg:1961:DBH

[Str61] Donald W. Strasburg. Diving behaviour of Hawaiian skipjack tuna. ICES Journal of Marine Science, 26(2):223–229, April 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/26/
2/223/651599.

Sundnes:1969:RDC

[SV69] Gunnar Sundnes and Eva Valen. Respiration of dry cysts of Artemia sauna L. ICES Journal of Marine Science, 32(3):413–415, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
32/3/413/714282.

Strasburg:1960:POT

[SY60] Donald W. Strasburg and Heeny S. H. Yuen. Progress in observing tuna underwater at sea. ICES Journal of Marine Science, 26 (1):80–93, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/26/1/80/2182129.

Symonds:1964:RSM

[Sym64] D. J. Symonds. Racial studies on Manx herring stocks. ICES Journal of Marine Science, 29(2):189–204, October 1964. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/29/2/189/
652939.

Szekielda:1968:CMA

[Sze68a] Karl-Heinz Szekielda. Ein Chemisches Modell für den Aufund Abbau Organischen Materials Und Dessen Anwendung in Der Offenen See. (German) [A chemical model for the build-up and breakdown of organic material and its application in the open sea]. ICES Journal of Marine Science, 32(2):180–187, November 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
2/180/756633.

Szekielda:1968:VUG

[Sze68b] Karl-Heinz Szekielda. Vergleichende Untersuchungen über den Gehalt an Organischem Kohlenstoff im Meerwasser Und Demkaliumpermanganatverbrauch. (German) [Comparative studies on the content of organic carbon in seawater and consumption of dempotassium permanganate]. ICES Journal of Marine Science, 32(1):17–24, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/32/1/17/641402.

Szekielda:1969:ACM

[Sze69a] Karl-Heinz Szekielda. The application of a chemical model to calculate the “theoretical” organic carbon in seawater. ICES Journal of Marine Science, 33(1):14–19, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/33/1/14/628743.

Szekielda:1969:DDC

[Sze69b] Karl-Heinz Szekielda. Le dosage du carbone particulaire dans l’eau de mer et son application dans le Golfe du Lion. (French) [The determination of particulate carbon in sea water and its application in the Gulf of Lion]. ICES Journal of Marine Science, 32(3):318–343, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
32/3/318/714272.

Taylor:1960:TGM

[Tay60a] Clyde C. Taylor. Temperature, growth, and mortality — the Pacific cockle. ICES Journal of Marine Science, 26(1):117–124, December 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
26/1/117/2182101.

Taylor:1960:WTC

[Tay60b] Clyde C. Taylor. Water temperature and cod growth-rate. ICES Journal of Marine Science, 25(2):223–224, March 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
2/223/623348.

Taylor:1962:GEM

[Tay62] Clyde C. Taylor. Growth equations with metabolic parameters. ICES Journal of Marine Science, 27(3):270–286, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
3/270/738927.

Thomas:1965:SVC

[TF65] H. J. Thomas and Maria José Figueiredo. Seasonal variations in the catch composition of the Norway lobster, Nephrops norvegicus (L.) around Scotland. ICES Journal of Marine Science, 30 (1):75–85, December 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/1/75/642396.

Teal:1962:TMR

[TH62] J. M. Teal and K. Halcrow. A technique for measurement of respiration of single copepods at sea. ICES Journal of Marine Science, 27(2):125–128, May 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/2/125/690277.

Talbot:1968:ARB

[TH68] J. W. Talbot and J. L. Henry. The adsorption of rhodamine-B on to materials carried in suspension by inshore waters. ICES Journal of Marine Science, 32(1):7–15, July 1968. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/32/1/7/641411.

Thomas:1964:SFN

[Tho64] H. J. Thomas. The spawning and fecundity of the Norway lobsters (Nephrops norvegicus L.) around the Scottish Coast. ICES Journal of Marine Science, 29(2):221–229, October 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/29/
2/221/652945.

Throndsen:1969:SMU

[Thr69] John Throndsen. A simple micropipette for use on the wild M 40 and the Zeiss plankton microscopes. ICES Journal of Marine Science, 32(3):430–432, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/32/3/430/714287.

Tungate:1965:IMD

[TM65] D. S. Tungate and D. Mummery. An inexpensive mechanical digital flowmeter. ICES Journal of Marine Science, 30(1):86, December 1965. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
30/1/86/642397.

Tranter:1960:MDZ

[Tra60] D. J. Tranter. A method for determining zooplankton volumes. ICES Journal of Marine Science, 25(3):272–278, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
3/272/666318.

Tungate:1967:NSC

[Tun67] D. S. Tungate. A new sedimentation chamber. ICES Journal of Marine Science, 31(2):284–285, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/31/2/284/696758.

Vaccaro:1962:OAS

[Vac62] R. F. Vaccaro. The oxidation of ammonia in sea water. ICES Journal of Marine Science, 27(1):3–14, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/27/1/3/581499.

Vatova:1961:PPH

[Vat61] Aristocle Vatova. Primary production in the High Venice Lagoon. ICES Journal of Marine Science, 26(2):148–155, April 1961. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/26/
2/148/651563.

Vatova:1962:SWF

[Vat62] Aristocle Vatova. The salt-water fish farms of the North Adriatic and their fauna. ICES Journal of Marine Science, 27(1): 109–114, January 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/27/1/109/581480.

Valdez:1962:ETS

[VC62] V. Valdez and D. H. Cushing. The effect of tidal streams on the presence of an extensive layer of midwater echo traces. ICES Journal of Marine Science, 27(3):236–247, November 1962. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/27/
3/236/738914.

Valdez:1966:DVD

[VC66] V. Valdez and D. H. Cushing. The diurnal variations in depth and quantity of echo traces and their distribution in area in the Southern Bight of the North Sea. ICES Journal of Marine Science, 30(2):237–254, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/30/2/237/702881.

Vaccaro:1960:MPD

[VR60] Ralph F. Vaccaro and John H. Ryther. Marine phytoplankton and the distribution of nitrite in the sea. ICES Journal of Marine Science, 25(3):260–271, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/25/3/260/666317.

Wood:1966:SDH

[WF66] R. J. Wood and M. Foster. Studies on the dunmore herring stock: 2. the otoliths of the dunmore herring. ICES Journal of Marine Science, 30(2):222–236, June 1966. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/30/2/222/702876.

Wilkins:1967:PWB

[Wil67a] N. P. Wilkins. Polymorphism of whole blood proteins in the cod (Gadus morhua L.). ICES Journal of Marine Science, 31(1):77–88, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
31/1/77/843996.

Williams:1967:MMH

[Wil67b] T. Williams. A method of measuring headless fish. ICES Journal of Marine Science, 31(2):279–283, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/31/2/279/696755.

Williams:1969:LHM

[Wil69a] C. S. Williams. The life history of Mytilicola intestinalis Steuer. ICES Journal of Marine Science, 32(3):419–429, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
3/419/714286.

Williams:1969:PVM

[Wil69b] C. S. Williams. Physical variations in Mytilicola intestinalis from two areas. ICES Journal of Marine Science, 32(3):435–437, April 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/32/
3/435/714292.

Waugh:1967:ERB

[WK67] G. Duncan Waugh and D. Key. Experiments with rhodamine B and European flat oysters (Ostrea edulis L). ICES Journal of Marine Science, 31(2):272–278, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/31/2/272/696751.

Wolfe:1967:LSG

[WS67] Douglas A. Wolfe and Claire L. Schelske. Liquid scintillation and Geiger counting efficiencies for carbon-14 incorporated by marine phytoplankton in productivity measurements. ICES Journal of Marine Science, 31(1):31–37, April 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/31/1/31/843967.

Zarnecki:1960:RCS

[Zar60] St. Zarnecki. Recent changes in the spawning habits of sea trout in the Upper Vistula. ICES Journal of Marine Science, 25(3):326–331, August 1960. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
25/3/326/666338.

Zarnecki:1964:MST

[ZDG64] S. Zarnecki, J. Duszyriski, and J. Gordziejczyk. Migrations of sea trout from Pomeranian rivers: Results of tagging in 1960. ICES Journal of Marine Science, 29(1):82–92, June 1964. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/29/1/82/598994.

Zijlstra:1963:RFP

[ZdV63] J. J. Zijlstra and J. F. de Veen. On the relative fishing power of Dutch trawlers. ICES Journal of Marine Science, 28(2): 241–245, September 1963. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/28/2/241/780217.

Zijlstra:1967:AEH

[Zij67] J. J. Zijlstra. Abundance estimates of herring in the North-Western North Sea. ICES Journal of Marine Science, 31(2):196–206, August 1967. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
31/2/196/696734.

Zijlstra:1969:RSN

[Zij69] J. J. Zijlstra. On the “racial” structure of North Sea autumnspawning herring. ICES Journal of Marine Science, 33(1):67–80, November 1969. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/33/1/67/628780.